### v2

# Ballistic Transport for Discrete Multi-Dimensional Schrödinger Operators With Decaying Potential

Abstract.
We consider the discrete Schrödinger operator $H=-\Delta+V$ on $\ell^{2}({\mathbb{Z}}^{d})$ with a decaying potential, in arbitrary lattice dimension $d\in{\mathbb{N}}^{*}$, where $\Delta$ is the standard discrete Laplacian and $V_{n}=o(|n|^{-1})$ as $|n|\to\infty$. We prove that the unitary evolution $e^{-{\rm i}tH}$ exhibits ballistic transport in the sense that, for any $r>0$, the weighted $\ell^{2}$-norm
$\|e^{-{\rm i}tH}u\|_{r}:=\left(\sum_{n\in{\mathbb{Z}}^{d}}(1+|n|^{2})^{r}|(e^{-{\rm i}tH}u)_{n}|^{2}\right)^{\frac{1}{2}}$
grows at rate $\simeq t^{r}$ as $t\to\infty$, provided that the initial state $u$ is in the absolutely continuous subspace and satisfies $\|u\|_{r}<\infty$.
The proof relies on commutator methods and a refined Mourre estimate, which yields quantitative lower bounds on transport for operators with purely absolutely continuous spectrum over appropriate spectral intervals. Compactness arguments and localized spectral projections are used to extend the result to perturbed operators, extending the classical result for the free Laplacian to a broader class of decaying potentials.
Key words and phrases:
Schrödinger operators, ballistic transport, Mourre estimate
D. D. was supported in part by NSF grants DMS–2054752 and DMS–2349919
1. Introduction and Main Result
1.1. General Setting and Questions
A main motivation to perform a spectral analysis of a Schrödinger operator $H=-\Delta+V$ is given by the known connections between spectral properties of $H$ and the long-term asymptotics of the associated unitary group $e^{-{\rm i}tH}$, with which one can describe the solutions of the associated time-dependent Schrödinger equation. That is, $\psi(t)=e^{-{\rm i}tH}\psi_{0}$ is the unique solution of the time-dependent equation
${\rm i}\partial_{t}\psi=H\psi,\qquad\left.\psi\right|_{t=0}=\psi_{0}.$
A fundamental result establishing such a connection is given by the RAGE theorem; compare, for example, [10, Theorem 1.6.7]. More refined connections can be found in [10, 23] and references therein.
Generally speaking, the more regular the spectral measure $\mu=\mu_{H,\psi_{0}}$ associated with the pair $(H,\psi_{0})$ is, the greater the tendency of $\psi$ is to leave compact regions in space as time grows.
In order to be more specific, let us consider the case where the operator $H$ acts in the Hilbert space $\ell^{2}({\mathbb{Z}}^{d})$; analogous statements exist in the case of the continuum, $L^{2}({\mathbb{R}}^{d})$, or the half line, $\ell^{2}({\mathbb{Z}}_{+})$ or $L^{2}({\mathbb{R}}_{+})$.
Consider the Lebesgue decomposition of $\mu$ into its absolutely continuous, singular continuous, and pure point parts:
$\mu=\mu_{\mathrm{ac}}+\mu_{\mathrm{sc}}+\mu_{\mathrm{pp}}.$
Let us first discuss the question of whether the time evolution leaves compact regions or is confined to a suitable compact region, up to an arbitrarily small error. The RAGE theorem states that
$\mu=\mu_{\mathrm{pp}}$ $\Longleftrightarrow$ $\forall\ \varepsilon>0,\ \exists\ N,\ \forall\ t\in{\mathbb{R}}\;:\;\sum_{|n|>N}|\langle\delta_{n},\psi(t)\rangle|^{2}<\varepsilon,$
$\mu=\mu_{\mathrm{sc}}+\mu_{\mathrm{ac}}$ $\Longleftrightarrow$ $\forall\ N\;:\;\lim_{T\to\infty}\frac{1}{2T}\int_{-T}^{T}\sum_{|n|\leq N}|\langle\delta_{n},\psi(t)\rangle|^{2}\,dt=0,$
$\mu=\mu_{\mathrm{ac}}$ $\Longrightarrow$ $\forall\ N\;:\;\lim_{|t|\to\infty}\sum_{|n|\leq N}|\langle\delta_{n},\psi(t)\rangle|^{2}=0.$
Recall that for a normalized initial state, $\|\psi_{0}\|=1$, $\sum_{|n|\leq N}|\langle\delta_{n},\psi(t)\rangle|^{2}$ gives the probability of the wavefunction being in the ball of radius $N$ centered at the origin at time $t$, so these statements indeed have the interpretation alluded to above.
We will be especially interested in cases where the spectral measure is absolutely continuous (a.c. for short). This is in response to the following general question: for initial states $\psi_{0}$ with $\mu=\mu_{\mathrm{ac}}$, is transport ballistic (formulated in Subsection 1.2)? There are general remarks that are relevant to this question (see the survey [13] for a much more in-depth discussion):
  • (i)
    One expects this to be true for many models of interest, but there is no general result.
  • (ii)
    In fact, there cannot be a completely general result as there are counterexamples, which can be constructed via inverse spectral theory in higher dimensions; see Bellissard and Schulz-Baldes [4].
Whether or not $\mu=\mu_{\mathrm{ac}}$ implies ballistic transport in one dimension being a difficult question in general, it is natural to explore it for specific interesting classes of potentials. This has indeed been done. For periodic $V$, it was shown by Asch and Knauf [3] in the continuum setting and Damanik, Lukic, and Yessen [12] in the discrete case. An extension to suitable limit-periodic potentials was obtained by Fillman [16] in the discrete case and by Young [31] in the continuum. Finally, suitable quasi-periodic potentials were studied by Zhao [33, 34], Zhang and Zhao [32], and Ge and Kachkovskiy [19]. In higher dimensions, there are results of Asch and Knauf [3], Black, Damanik, Malinovitch, and Young [6], Boutet de Monvel and Sabri [7], Fillman [17], and Karpeshina, Lee, Parnovski, Shterenberg, and Stolz [21, 22]. [^1] It should be noted, however, that [21, 22] consider the time-averaged evolution, whereas this paper and the other papers mentioned above do not.
While periodic, limit-periodic, and quasi-periodic potentials are of course of great importance with respect to actual physical models, another very important class is that of decaying potentials, and the absence of a ballistic transport result for them is quite striking. The present paper seeks to address this and produce a ballistic transport result for suitable decaying potentials.
1.2. Formulation and Main Result
For $r\in{\mathbb{R}}_{+}$ and $d\in{\mathbb{N}}^{*}$, let us define the subspace ${\mathcal{H}}^{r}({\mathbb{Z}}^{d})$ of $\ell^{2}({\mathbb{Z}}^{d})$, as
${\mathcal{H}}^{r}({\mathbb{Z}}^{d}):=\Big\{u=(u_{n})\in\ell^{2}({\mathbb{Z}}^{d}):\sum_{n\in{\mathbb{Z}}^{d}}(1+|n|^{2})^{r}|u_{n}|^{2}<\infty\Big\},$
with $|n|:=\sqrt{n_{1}^{2}+\cdots+n_{d}^{2}}$. It also naturally defines the weighted $\ell^{2}$-norm $\|\cdot\|_{r}$ on ${\mathcal{H}}^{r}({\mathbb{Z}}^{d})$ via
$\|u\|_{r}:=\left(\sum_{n\in{\mathbb{Z}}^{d}}(1+|n|^{2})^{r}|u_{n}|^{2}\right)^{1/2}.$
We identify the $\|\cdot\|_{\ell^{2}}$-norm with the $\|\cdot\|_{0}$-norm.
We consider the $d$-dimensional discrete Schrödinger operator $H=-\Delta+V$ defined through the discrete Laplacian
(1) $\Delta:\ell^{2}({\mathbb{Z}}^{d})\to\ell^{2}({\mathbb{Z}}^{d}),\quad(\Delta u)_{n}=\sum_{j=1}^{d}(\Delta_{j}u)_{n}:=\sum_{j=1}^{d}(u_{n+e_{j}}+u_{n-e_{j}}),$
with $\{e_{j}\}_{j=1}^{d}$ the canonical basis of ${\mathbb{Z}}^{d}$, and the multiplication by the potential $V=(V_{n})$.
Let ${\mathcal{H}}_{\rm ac}(H)$ be the absolutely continuous spectral subspace associated with $H$. For $r>0$, let ${\mathcal{H}}_{\rm ac}^{r}({\mathbb{Z}}^{d}):={\mathcal{H}}^{r}({\mathbb{Z}}^{d})\cap{\mathcal{H}}_{\rm ac}(H)$. Here is our main result:
Theorem 1.1.
If $V=(V_{n})$ satisfies
(2) $V_{n}=o(|n|^{-1}),\quad|n|\to\infty,$
then, for the Schrödinger operator $H=-\Delta+V$, $e^{-{\rm i}tH}$ exhibits ballistic transport in the sense that, for any $u\in{\mathcal{H}}_{\rm ac}^{r}({\mathbb{Z}}^{d})\setminus\{0\}$ and any $r>0$, there exists a constant $C_{u,r}>1$ such that
(3) $C_{u,r}^{-1}\leq\frac{\|e^{-{\rm i}tH}u\|_{r}}{t^{r}}\leq C_{u,r}\quad\forall\ t\geq 1.$
Remark 1.2.
In the terminology of the survey [13], Theorem 1.1 establishes ballistic transport in the norm sense for all moments under the decay assumption (2).
The condition (2) arises naturally from our proof. In addition, it is also known to be the sharp decay condition that ensures that the interior of the essential spectrum is purely a.c. in one dimension [30]. Note that Wigner-von Neumann potentials satisfy $V_{n}=O(|n|^{-1})$, $|n|\to\infty$, and allow for eigenvalues embedded in the interior of the essential spectrum [11, 29]. [^2] See especially [11, Example on p.52] for how to embed an eigenvalue in the discrete case. Whether or not (2) ensures purely a.c. interior essential spectrum in higher dimensions as well is not known (to the best of our knowledge). We will show the absence of singular continuous spectrum in Proposition 3.2 below, but at this point we cannot rule out the existence of isolated eigenvalues embedded in the interior of the essential spectrum. Clearly, the ballistic lower bound in (3) cannot hold for eigenstates, and hence we naturally restrict our attention in Theorem 1.1 to initial states in the absolutely continuous subspace.
2. Mourre Estimate and Order1-1 Ballistic Lower Bound
Define the weight operator $Q:{\mathcal{H}}^{1}({\mathbb{Z}}^{d})\to\ell^{2}({\mathbb{Z}}^{d})$ by
$(Qu)_{n}:=\sqrt{1+|n|^{2}}\,u_{n},\quad n\in{\mathbb{Z}}^{d}.$
2.1. Spectral Projection and Weight Operator
For a self-adjoint operator $H$ on $\ell^{2}({\mathbb{Z}}^{d})$, and $I\subset{\mathbb{R}}$, let $\chi_{I}(H)$ be the spectral projection of $H$ corresponding to $I$.
Lemma 2.1.
If $[Q,H]$ can be extended as a bounded operator on $\ell^{2}({\mathbb{Z}}^{d})$, then $\chi_{I}(H)u\in{\mathcal{H}}^{1}({\mathbb{Z}}^{d})$ for any $I\subset{\mathbb{R}}$ and any $u\in{\mathcal{H}}^{1}({\mathbb{Z}}^{d})$.
Proof.
Since $[Q,H]$ can be extended as a bounded operator on $\ell^{2}({\mathbb{Z}}^{d})$, for any $z\in{\mathbb{C}}\setminus{\mathbb{R}}$, it holds on $\ell^{2}({\mathbb{Z}}^{d})$ that
$Q,(H-z)^{-1}]=-(H-z)^{-1}[Q,H](H-z)^{-1}.$
Through the Helffer-Sjöstrand representation [20], for $\varphi\in C_{c}^{\infty}({\mathbb{R}})$, we have that $Q,\varphi(H)]$ can be extended as a bounded operator on $\ell^{2}({\mathbb{Z}}^{d})$.
Now, for $\chi_{I}$, there exists a sequence $\{\varphi_{n}\}\subset C_{c}^{\infty}({\mathbb{R}})$ such that $\sup|\varphi_{n}|\leq 1$ and $\varphi_{n}\to\chi_{I}$ pointwise. By the uniform bound $\|[Q,\varphi_{n}(H)]\|_{{\mathcal{B}}(\ell^{2}({\mathbb{Z}}^{d}))}\leq C$ for some $C>0$, as well as the strong convergence $Q,\varphi_{n}(H)]\to[Q,\chi_{I}(H)]$, we have $\|[Q,\chi_{I}(H)]\|_{{\mathcal{B}}(\ell^{2}({\mathbb{Z}}^{d}))}\leq C$. Since $\chi_{I}(H)$ is bounded on $\ell^{2}({\mathbb{Z}}^{d})$,
$Q\chi_{I}(H)u=\chi_{I}(H)Qu+[Q,\chi_{I}(H)]u\in\ell^{2}({\mathbb{Z}}^{d}),\quad\forall\ u\in{\mathcal{H}}^{1}({\mathbb{Z}}^{d}).\qed$
Proposition 2.2.
Let $H$ be a self-adjoint operator on $\ell^{2}({\mathbb{Z}}^{d})$ with non-empty a.c. spectrum. Assume that $[Q,H]$ can be extended as a bounded operator on $\ell^{2}({\mathbb{Z}}^{d})$. Let $u\in{\mathcal{H}}_{\rm ac}^{1}({\mathbb{Z}}^{d})$ with
(4) $e^{-{\rm i}tH}u\in{\mathcal{H}}^{1}({\mathbb{Z}}^{d}),\quad\forall\ t\in{\mathbb{R}}.$
Given $I,J\subset{\mathbb{R}}$ with $0<\mathrm{Leb}(I),\ \mathrm{Leb}(J)<\infty$ and $I\cap J=\emptyset$, we have, for $u\in{\mathcal{H}}_{\rm ac}^{1}({\mathbb{Z}}^{d})$,
(5) $Qe^{-{\rm i}tH}\chi_{I}(H)u,\ Qe^{-{\rm i}tH}\chi_{J}(H)u\in\ell^{2}({\mathbb{Z}}^{d}),\quad\forall\ t\in{\mathbb{R}},$
and for $t>0$ sufficiently large,
(6) $\left\|Qe^{-{\rm i}tH}(\chi_{I}(H)u+\chi_{J}(H)u)\right\|_{0}$
$>$ $\frac{1}{2}\max\left\{\|Qe^{-{\rm i}tH}\chi_{I}(H)u\|_{0},\ \|Qe^{-{\rm i}tH}\chi_{J}(H)u\|_{0}\right\}.$
Remark 2.3.
The assumption (4) holds true for every $u\in{\mathcal{H}}^{1}({\mathbb{Z}}^{d})$ if the time evolution $e^{-{\rm i}tH}$ of the operator $H$ exhibits a ballistic upper bound (cf. (16) in Proposition 3.1). As such a ballistic upper bound is established for Schrödinger operators in Proposition 3.1, these operators satisfy the assumption (4).
Proof.
Since $e^{-{\rm i}tH}u\in{\mathcal{H}}^{1}({\mathbb{Z}}^{d})$ for every $t\in{\mathbb{R}}$, we obtain (5) through Lemma 2.1, by noting that
$e^{-{\rm i}tH}\chi_{I}(H)u=\chi_{I}(H)e^{-{\rm i}tH}u,\quad e^{-{\rm i}tH}\chi_{J}(H)u=\chi_{J}(H)e^{-{\rm i}tH}u.$
Direct computations yield that
$\|Qe^{-{\rm i}tH}(\chi_{I}(H)u+\chi_{J}(H)u)\|^{2}_{0}$
$=$ $\|Qe^{-{\rm i}tH}\chi_{I}(H)u\|^{2}_{0}+\|Qe^{-{\rm i}tH}\chi_{J}(H)u\|^{2}_{0}+C^{u}_{I,J}(t)+C^{u}_{J,I}(t),$
with $C^{u}_{I,J}(t)$ and $C^{u}_{J,I}(t)$ defined as
$C^{u}_{I,J}(t)$ $:=$ $\left\langle Qe^{-{\rm i}tH}\chi_{I}(H)u,Qe^{-{\rm i}tH}\chi_{J}(H)u\right\rangle,$
$C^{u}_{J,I}(t)$ $:=$ $\left\langle Qe^{-{\rm i}tH}\chi_{J}(H)u,Qe^{-{\rm i}tH}\chi_{I}(H)u\right\rangle.$
In view of (5), in order to show (6), it is sufficient to show that $C^{u}_{I,J}(t)$, $C^{u}_{J,I}(t)\to 0$ as $t\to\infty$. Let $E_{H}(\cdot)$ be the projection-valued measure of $H$ and let $\{\delta_{n}\}_{n\in{\mathbb{Z}}^{d}}$ be the standard basis of $\ell^{2}({\mathbb{Z}}^{d})$. Define the scalar measure $\mu_{n}(\cdot)=\langle\delta_{n},E_{H}(\cdot)u\rangle$, which is a.c. on $\sigma_{\rm ac}(H)$ since $u\in{\mathcal{H}}_{\rm ac}^{1}({\mathbb{Z}}^{d})$. Then we have
$(e^{-{\rm i}tH}\chi_{I}(H)u)_{n}=\int_{I}e^{-{\rm i}t\lambda}\,d\mu_{n}(\lambda),\quad(e^{-{\rm i}tH}\chi_{J}(H)u)_{n}=\int_{J}e^{-{\rm i}t\sigma}\,d\mu_{n}(\sigma),$
and hence, by Fubini’s theorem,
$C^{u}_{I,J}(t)$ $=$ $\sum_{n\in{\mathbb{Z}}^{d}}(1+|n|^{2})\left(\int_{I}e^{-{\rm i}t\lambda}\,d\mu_{n}(\lambda)\right)\overline{\left(\int_{J}e^{-{\rm i}t\sigma}\,d\mu_{n}(\sigma)\right)}$
$=$ $\int_{I}\int_{J}e^{-{\rm i}t(\lambda-\sigma)}\sum_{n\in{\mathbb{Z}}^{d}}(1+|n|^{2})\,d\mu_{n}(\lambda)\,d\bar{\mu}_{n}(\sigma)$
$=:$ $\int_{{\mathbb{R}}}e^{-{\rm i}tE}\,d\nu(E),$
where the measure $\nu$ is defined as: for $B\subset{\mathbb{R}}$,
$\nu(B)=\sum_{n\in{\mathbb{Z}}^{d}}(1+|n|^{2})\int_{I\times J}\chi_{\{\lambda-\sigma\in B\}}(\lambda,\sigma)\,d\mu_{n}(\lambda)\,d\bar{\mu}_{n}(\sigma).$
Since $u\in{\mathcal{H}}_{\rm ac}^{1}({\mathbb{Z}}^{d})$, we have
$|\nu({\mathbb{R}})|$ $\leq$ $\sum_{n\in{\mathbb{Z}}^{d}}(1+|n|^{2})\left|\mu_{n}(I)\right|\left|\bar{\mu}_{n}(J)\right|$
$\leq$ $\sum_{n\in{\mathbb{Z}}^{d}}(1+|n|^{2})|(\chi_{I}(H)u)_{n}||(\chi_{J}(H)\bar{u})_{n}|$
$\leq$ $\frac{1}{2}(\|\chi_{I}(H)u\|_{1}^{2}+\|\chi_{J}(H)u\|_{1}^{2}).$
According to the Riemann-Lebesgue lemma, $C^{u}_{I,J}(t)\to 0$ as $t\to\infty$. Similarly, $C^{u}_{J,I}(t)\to 0$.∎
2.2. Strict Mourre Estimate
From now on, for the self-adjoint operator $H$, we always assume that $[Q,H]$ can be extended to a bounded operator on $\ell^{2}({\mathbb{Z}}^{d})$, such that Lemma 2.1 and Proposition 2.2 are applicable. We assume further that $Q^{2},H]$ can be extended to a bounded operator from ${\mathcal{H}}^{1}({\mathbb{Z}}^{d})$ to $\ell^{2}({\mathbb{Z}}^{d})$. We will verify these assumptions for the class of Schrödinger operators of interest to us in Example 2.7 below.
Proposition 2.4.
For a self-adjoint operator $H$ acting on $\ell^{2}({\mathbb{Z}}^{d})$, assume that, on $I\subset{\mathbb{R}}$ with $\mathrm{Leb}(I\cap\sigma(H))>0$, there is $\theta>0$ such that
(7) $\chi_{I}(H)[H,[H,-Q^{2}]]\chi_{I}(H)\geq\theta\chi_{I}(H).$
Then, for any $u\in{\mathcal{H}}_{\rm ac}^{1}({\mathbb{Z}}^{d})$ satisfying (4), we have
(8) $\frac{\|e^{-{\rm i}tH}u\|_{1}}{t}\gtrsim\theta^{\frac{1}{2}}\|\chi_{I}(H)u\|_{0},\quad\forall\ t>0.$
Remark 2.5.
(a) It is well known that the strict Mourre estimate implies purely a.c. spectrum; compare [2, Section 7.2]. Hence, under the assumption (7), $\chi_{I}(H)u\in{\mathcal{H}}_{\rm ac}(H)$ for any $u\in\ell^{2}({\mathbb{Z}}^{d})$.
(b) Beyond the $\ell^{2}({\mathbb{Z}}^{d})$ setting, the Mourre estimate has also been employed to establish the absence of singular continuous spectrum for adjacency operators on graphs [24]. Moreover, it plays an important role in the study of scattering theory; see, e.g., [5, 15, 18].
The proof of Proposition 2.4 relies on the following lemma.
Lemma 2.6.
Under the assumption of Proposition 2.4, we have that
(9) $\frac{\|Qe^{-{\rm i}tH}\chi_{I}(H)u\|_{0}}{t}\gtrsim\theta^{\frac{1}{2}}\|\chi_{I}(H)u\|_{0},\quad\forall\ t>0.$
Proof.
Based on the proofs of [9, Proposition 1.2] and [13, Proposition 4.6], we have
(10) $\|Qe^{-{\rm i}tH}\chi_{I}(H)u\|_{0}^{2}$ $=$ $\langle\chi_{I}(H)u,e^{{\rm i}tH}Q^{2}e^{-{\rm i}tH}\chi_{I}(H)u\rangle$
$=$ $\|Q\chi_{I}(H)u\|_{0}^{2}-{\rm i}t\langle\chi_{I}(H)u,[H,-Q^{2}]\chi_{I}(H)u\rangle$
$+\,\int_{0}^{t}\int_{0}^{s}\left\langle e^{-{\rm i}\sigma H}\chi_{I}(H)u,[H,[H,-Q^{2}]]e^{-{\rm i}\sigma H}\chi_{I}(H)u\right\rangle\,d\sigma\,ds,$
where the assumption $\chi_{I}(H)u\in{\mathcal{H}}^{1}({\mathbb{Z}}^{d})$ in [13, Proposition 4.6] is guaranteed by Lemma 2.1 once we have $u\in{\mathcal{H}}^{1}({\mathbb{Z}}^{d})$. Each term in the above computation is finite for any finite $t$, under the assumption that $Q^{2},H]$ is bounded on ${\mathcal{H}}^{1}({\mathbb{Z}}^{d})$.
As $t\to\infty$, we obtain (9), since the strict Mourre estimate (7) implies that the integral term in (10) satisfies
$\left|\int_{0}^{t}\int_{0}^{s}\left\langle e^{-{\rm i}\sigma H}\chi_{I}(H)u,[H,[H,-Q^{2}]]e^{-{\rm i}\sigma H}\chi_{I}(H)u\right\rangle\,d\sigma\,ds\right|\geq\frac{\theta\|\chi_{I}(H)u\|_{0}^{2}}{2}t^{2}.\qed$
Proof of Proposition 2.4. Write $\sigma(H)=I\cup I^{c}$ with $I$ and $I^{c}$ mutually disjoint. For $u\in{\mathcal{H}}_{\rm ac}^{1}({\mathbb{Z}}^{d})$ satisfying (4), we have
$u=\chi_{I}(H)u+\chi_{I^{c}}(H)u,$
and according to Proposition 2.2, we have
$\|Qe^{-{\rm i}tH}u\|_{0}\geq\frac{1}{2}\|Qe^{-{\rm i}tH}\chi_{I}(H)u\|_{0}$
for $t>0$ sufficiently large, which together with (9) implies (8).∎
Example 2.7.
Let us apply Proposition 2.4 to the free Schrödinger case $H=-\Delta$ with the $d$-dimensional Laplacian operator defined as in (1), whose time evolution is well known to exhibit ballistic transport. Note that for $u\in{\mathcal{H}}^{1}({\mathbb{Z}}^{d})$, we have
$([Q,-\Delta]u)_{n}$ $=$ $\sum_{j=1}^{d}\left(\sqrt{1+|n+e_{j}|^{2}}-\sqrt{1+|n|^{2}}\right)u_{n+e_{j}}$
$+\,\sum_{j=1}^{d}\left(\sqrt{1+|n-e_{j}|^{2}}-\sqrt{1+|n|^{2}}\right)u_{n-e_{j}},$
where, for $1\leq j\leq d$,
$\left|\sqrt{1+|n+e_{j}|^{2}}-\sqrt{1+|n|^{2}}\right|$ $=$ $\sqrt{1+|n|^{2}}\left|\left(1+\frac{2n_{j}+1}{1+|n|^{2}}\right)^{\frac{1}{2}}-1\right|$
$\leq$ $\frac{2|n_{j}|+1}{\sqrt{1+|n|^{2}}}\ \leq\ \sqrt{5},$
$\left|\sqrt{1+|n-e_{j}|^{2}}-\sqrt{1+|n|^{2}}\right|$ $\leq$ $\sqrt{5}.$
Hence, $Q,-\Delta]$ can be extended to a bounded operator on $\ell^{2}({\mathbb{Z}}^{d})$. Moreover, for $u\in{\mathcal{H}}^{1}({\mathbb{Z}}^{d})$,
$([Q^{2},-\Delta]u)_{n}$ $=$ $\sum_{j=1}^{d}\left(|n+e_{j}|^{2}-|n|^{2}\right)u_{n+e_{j}}$
$+\sum_{j=1}^{d}\left(|n-e_{j}|^{2}-|n|^{2}\right)u_{n-e_{j}}$
$=$ $\sum_{j=1}^{d}\left(\left(2n_{j}+1\right)u_{n+e_{j}}-\left(2n_{j}-1\right)u_{n-e_{j}}\right),$
which shows that $Q^{2},-\Delta]$ can be extended to a bounded operator on ${\mathcal{H}}^{1}({\mathbb{Z}}^{d})$. Since any potential $V$ commutes with $Q$ and $Q^{2}$, it follows from the considerations above that for every discrete Schrödinger operator $H=-\Delta+V$, $[Q,H]$ and $Q^{2},H]$ can be extended to be bounded on $\ell^{2}({\mathbb{Z}}^{d})$ and ${\mathcal{H}}^{1}({\mathbb{Z}}^{d})$, respectively.
By direct computations, we have
(11) $\Delta,[-\Delta,-Q^{2}]]=2\sum_{j=1}^{d}(4{\rm I}-\Delta_{j}^{2}).$
Given $0<\theta<2d$, on the sub-interval
(12) $J_{\theta}:=]-2d+\theta,2d-\theta[\,\subset[-2d,2d]=\sigma(-\Delta),$
we have that
$\chi_{J_{\theta}}(-\Delta)[-\Delta,[-\Delta,-Q^{2}]]\chi_{J_{\theta}}(-\Delta)\geq\frac{2}{d}\theta\,\chi_{J_{\theta}}(-\Delta).$
Since for any $u\in{\mathcal{H}}^{1}({\mathbb{Z}}^{d})\setminus\{0\}$, there exists $\theta>0$ such that $\|\chi_{J_{\theta}}(-\Delta)u\|_{0}>0$, through Proposition 2.4, there exists a constant $c>0$, independent of $t$, such that
$\frac{\|e^{{\rm i}t\Delta}u\|_{1}}{t}>c,\quad\forall\ t>0.$
2.3. Mourre Estimate With a Compact Perturbation
According to Proposition 2.4, the strict Mourre estimate (7), which implies purely a.c. spectrum of $H$ on $I$ [2, Section 7.2], also implies ballistic transport for any initial state with non-trivial projection onto ${\rm Ran}\,\chi_{I}(H)$. This can be suitably generalized for the general Mourre estimate on $I$: there exist $\theta>0$ and a self-adjoint compact operator $K$ such that
(13) $\chi_{I}(H)[H,[H,-Q^{2}]]\chi_{I}(H)\geq\theta\chi_{I}(H)+K.$
In view of [26] or [2, Corollary 7.2.11], the estimate (13) implies that in $I$, the singular continuous spectrum of $H$ is empty and all eigenvalues of $H$ are isolated.
Proposition 2.8.
For a self-adjoint operator $H$ acting on $\ell^{2}({\mathbb{Z}}^{d})$, assume that on $I\subset{\mathbb{R}}$ with $\mathrm{Leb}(I\cap\sigma(H))>0$, the estimate (13) holds. Then there exists a subset $I_{1}\subset I$ with $\mathrm{Leb}(I_{1}\cap\sigma(H))>0$ such that
(14) $\chi_{I_{1}}(H)[H,[H,-Q^{2}]]\chi_{I_{1}}(H)\geq\frac{\theta}{2}\chi_{I_{1}}(H).$
Proof.
Similar to the proof of Lemma 2.13 in [25], we show that, by shrinking $I$ enough, a strict Mourre estimate (14) is true.
Take $E_{0}\in I\cap\left(\sigma_{ac}\left(H\right)\setminus\sigma_{pp}\left(H\right)\right)$ and $\delta>0$, define the subset $I_{1}(\delta)$ as
$I_{1}(\delta):=\left]E_{0}-\delta,E_{0}+\delta\right[\cap I\quad{\rm s.\ t.}\quad I_{1}(\delta)\cap\sigma_{pp}\left(H\right)=\emptyset.$
This is possible since the eigenvalues of $H$ in $I$ are isolated. If $\delta$ is sufficiently small, we have in addition
(15) $\left\|\chi_{I_{1}(\delta)}\left(H\right)K\chi_{I_{1}(\delta)}\left(H\right)\right\|_{{\mathcal{H}}^{0}\to{\mathcal{H}}^{0}}\leq\frac{\theta}{2}.$
Indeed, all spectral measures of $H$ are purely a.c. on $I_{1}(\delta)$, which means that
$\varphi_{\delta}:=\chi_{I_{1}(\delta)}\left(H\right)\varphi\in{\mathcal{H}}_{\rm ac}(H),\qquad\forall\ \varphi\in{\mathcal{H}}^{0}({\mathbb{Z}}^{d}).$
Then, by the dominated convergence theorem, we have
$\left\|\chi_{I_{1}(\delta)}\left(H\right)\varphi\right\|_{0}^{2}=\int_{\mathbb{R}}\chi_{I_{1}(\delta)}\left(E\right)^{2}\,d\mu_{H,\varphi}(E)\rightarrow 0\quad{\rm as}\quad\delta\rightarrow 0.$
In particular, $\chi_{I_{1}(\delta)}\left(H\right)\rightarrow 0$ strongly as $\delta\rightarrow 0$. Since $K$ is compact, $\chi_{I_{1}(\delta)}\left(H\right)K\rightarrow 0$ uniformly as $\delta\rightarrow 0$. Therefore, for $\delta$ sufficiently small, (15) holds true.
Combining this with (13), we have that
$\chi_{I_{1}(\delta)}\left(H\right)[H,[H,-Q^{2}]]\chi_{I_{1}(\delta)}\left(H\right)$
$=$ $\chi_{I_{1}(\delta)}\left(H\right)\chi_{I}\left(H\right)[H,[H,-Q^{2}]]\chi_{I}\left(H\right)\chi_{I_{1}(\delta)}\left(H\right)$
$\geq$ $\theta\chi_{I_{1}(\delta)}\left(H\right)\chi_{I}\left(H\right)\chi_{I_{1}(\delta)}\left(H\right)+\chi_{I_{1}(\delta)}\left(H\right)K\chi_{I_{1}(\delta)}\left(H\right)$
$\geq$ $\frac{\theta}{2}\chi_{I_{1}(\delta)}\left(H\right),$
which implies (14).∎
Under the assumption (13) on $I$, for any non-vanishing $u\in{\mathcal{H}}_{\rm ac}(H)$, there is one subset $I_{1}$ such that $\|\chi_{I_{1}}(H)u\|_{0}>0$. Then, according to Proposition 2.4, we have
Corollary 2.9.
For a self-adjoint operator $H$ on $\ell^{2}({\mathbb{Z}}^{d})$, if it satisfies the Mourre estimate (13), then $e^{-{\rm i}tH}$ exhibits the order1-1 ballistic lower bound: $\forall\ u\in{\mathcal{H}}_{\rm ac}^{1}({\mathbb{Z}}^{d})$, there exists $I_{1}\subset I$ with ${\rm Leb}(I_{1}\cap\sigma(H))>0$ such that
$\frac{\|e^{-{\rm i}tH}u\|_{1}}{t}\gtrsim\theta^{\frac{1}{2}}\|\chi_{I_{1}}(H)u\|_{0},\quad\forall\ t>0.$
3. Proof of Theorem 1.1
In this section, we provide a proof of Theorem 1.1. We will split it into two parts, presented in separate subsections, one addressing the upper bound and one addressing the lower bound in (3).
3.1. Ballistic Upper Bound
By general principles, the upper bound in (3) holds for any $r>0$ and for any self-adjoint discrete Schrödinger operator. That is, all we require is that the potential is real-valued, $V:{\mathbb{Z}}^{d}\to{\mathbb{R}}$, in which case the associated Schrödinger operator $H=-\Delta+V$ is self-adjoint on the domain
$D(H)=D(V)=\{u\in\ell^{2}({\mathbb{Z}}^{d}):Vu\in\ell^{2}({\mathbb{Z}}^{d})\}.$
Such a result is well known; compare, for example, the discussions in [1, Appendix B], [6, Lemma B.2], [7, Theorem A.1], [10, Theorem 2.6.2], [14, Theorem 2.22], and [32, Theorem 3]. [^3] There is also a ballistic upper bound result in the continuum case due to Radin and Simon [27]. For the reader’s convenience, we include a proof below. In fact, in doing so we establish a particular instance of the upper bound, (17), that is crucial in our discussion of the orderr-r ballistic lower bound for $0<r<1$ below; compare Lemma 3.5.
Proposition 3.1.
For any self-adjoint discrete Schrödinger operator $H=-\Delta+V$ on $\ell^{2}({\mathbb{Z}}^{d})$, we have the ballistic upper bound: there exists a constant $c_{r}>0$, independent of $u$, such that
(16) $\|e^{-{\rm i}tH}u\|_{r}\leq c_{r}\|u\|_{r}t^{r},\quad\forall\ u\in{\mathcal{H}}^{r}({\mathbb{Z}}^{d}),\;t\geq 1.$
Moreover, there exists $\tilde{c}_{2}>0$ such that
(17) $\|e^{-{\rm i}tH}u\|_{2}\leq\|u\|_{2}+\|u\|_{1}t+\tilde{c}_{2}\|u\|_{0}t^{2},\quad\forall\ u\in{\mathcal{H}}^{2}({\mathbb{Z}}^{d}),\;t\geq 1.$
Proof.
Given $m\in{\mathbb{N}}^{*}$, for $u\in{\mathcal{H}}^{m}({\mathbb{Z}}^{d})$, we have
(18) $e^{{\rm i}tH}Q^{m}e^{-{\rm i}tH}u=Q^{m}u+{\rm i}\int_{0}^{t}e^{{\rm i}sH}[H,Q^{m}]e^{-{\rm i}sH}u\,ds,$
where $H,Q^{m}]$ is well-defined on ${\mathcal{H}}^{m-1}({\mathbb{Z}}^{d})$. Indeed, noting that $H,Q^{m}]=[Q^{m},\Delta]$ for $H=-\Delta+V$ for any potential $V$, we have, for $\psi\in{\mathcal{H}}^{m-1}({\mathbb{Z}}^{d})$ and $n\in{\mathbb{Z}}^{d}$,
$([H,Q^{m}]\psi)_{n}$ $=$ $([Q^{m},\Delta]\psi)_{n}$
$=$ $\sum_{j=1}^{d}\left((1+|n|^{2})^{\frac{m}{2}}-(1+|n+e_{j}|^{2})^{\frac{m}{2}}\right)\psi_{n+e_{j}}$
$+\,\sum_{j=1}^{d}\left((1+|n|^{2})^{\frac{m}{2}}-(1+|n-e_{j}|^{2})^{\frac{m}{2}}\right)\psi_{n-e_{j}},$
which implies that
$\|[H,Q^{m}]\psi\|_{0}\lesssim_{m}\|Q^{m-1}\psi\|_{0}=\|\psi\|_{m-1},$
since, for $j=1,\cdots,d$, for $n\in{\mathbb{Z}}^{d}$,
$\left|(1+|n|^{2})^{\frac{m}{2}}-(1+|n+e_{j}|^{2})^{\frac{m}{2}}\right|$
$\leq$ $\sum_{k=1}^{m}\binom{m}{k}\left|\langle n+e_{j},n+e_{j}\rangle^{\frac{k}{2}}-\langle n,n\rangle^{\frac{k}{2}}\right|$
$\leq$ $\sum_{k=1}^{m}\binom{m}{k}\sum_{l=1}^{k}\binom{k}{l}|2n_{j}+1|^{\frac{l}{2}}\langle n+e_{j},n+e_{j}\rangle^{\frac{k-l}{2}}$
$\lesssim_{m}$ $(1+|n+e_{j}|^{2})^{\frac{m-1}{2}},$
and similarly,
$\left|(1+|n|^{2})^{\frac{m}{2}}-(1+|n-e_{j}|^{2})^{\frac{m}{2}}\right|\lesssim_{m}(1+|n-e_{j}|^{2})^{\frac{m-1}{2}}.$
Hence, when $V$ is a real potential, we have, through (18),
$\|e^{-{\rm i}tH}u\|_{m}$ $=$ $\|e^{{\rm i}tH}Q^{m}e^{-{\rm i}tH}u\|_{0}$
$\leq$ $\|Q^{m}u\|_{0}+\int_{0}^{t}\|[H,Q^{m}]e^{-{\rm i}sH}u\|_{0}\,ds$
$\lesssim_{m}$ $\|u\|_{m}+\int_{0}^{t}\|e^{-{\rm i}sH}u\|_{m-1}\,ds,\quad\forall\ t\geq 0.$
Since for $m=1$, $\|e^{-{\rm i}tH}u\|_{m-1}=\|u\|_{0}$, we obtain the linear upper bound for $\|e^{-{\rm i}tH}u\|_{1}$ for $u\in{\mathcal{H}}^{1}({\mathbb{Z}}^{d})$: there exists $\tilde{c}_{1}>0$ such that
$\|e^{-{\rm i}tH}u\|_{1}\leq\|u\|_{1}+\tilde{c}_{1}\|u\|_{0}t,\quad\forall\ t\geq 0.$
Then, via induction, for $u\in{\mathcal{H}}^{2}({\mathbb{Z}}^{d})$,
$\|e^{-{\rm i}tH}u\|_{2}$ $\leq$ $\|u\|_{2}+\int_{0}^{t}\|e^{-{\rm i}sH}u\|_{1}\,ds$
$\leq$ $\|u\|_{2}+\|u\|_{1}t+\frac{\tilde{c}_{1}}{2}\|u\|_{0}t^{2},\quad\forall\ t\geq 0.$
which shows (17). The above inequalities also imply that, for $m=1,2$,
$\|e^{-{\rm i}tH}u\|_{m}\leq c^{\prime}_{m}\|u\|_{m}(1+t)^{m},\quad\forall\ t\geq 0,$
through which we obtain (16) for $r=1,2$.
Assume that, for $m\in{\mathbb{N}}^{*}$, there exists a constant $c^{\prime}_{m}>0$ such that, for any $u\in{\mathcal{H}}^{m}({\mathbb{Z}}^{d})$,
$\|e^{-{\rm i}tH}u\|_{m}\leq c^{\prime}_{m}\|u\|_{m}(1+t)^{m},\quad\forall\ t\geq 0.$
By induction, we have
$\|e^{-{\rm i}tH}u\|_{m+1}$ $\lesssim_{m}$ $\|u\|_{m+1}+\int_{0}^{t}\|e^{-{\rm i}sH}u\|_{m}\,ds$
$\lesssim_{m}$ $\|u\|_{m+1}+\frac{c^{\prime}_{m}}{m+1}\|u\|_{m}(1+t)^{m+1},\quad\forall\ t\geq 0.$
Therefore, the upper bound (16) is shown when $r>0$ is an integer.
For $r\in\,]m,m+1[$ with some $m\in{\mathbb{N}}$ ($\{0,1,2,\ldots\}$) we have
$r=\left((m+1)-r\right)m+\left(r-m\right)(m+1).$
Applying Hölder’s inequality, we obtain, for $\psi\in{\mathcal{H}}^{m+1}({\mathbb{Z}}^{d})$,
(19) $\|\psi\|^{2}_{r}$ $=$ $\sum_{n\in{\mathbb{Z}}^{d}}(1+|n|^{2})^{\left((m+1)-r\right)m+\left(r-m\right)(m+1)}\left|\psi_{n}\right|^{2}$
$\leq$ $\left(\sum_{n\in{\mathbb{Z}}^{d}}(1+|n|^{2})^{m}\left|\psi_{n}\right|^{2}\right)^{(m+1)-r}\left(\sum_{n\in{\mathbb{Z}}^{d}}(1+|n|^{2})^{m+1}\left|\psi_{n}\right|^{2}\right)^{r-m}$
$=$ $\|\psi\|_{m}^{2(m+1-r)}\|\psi\|_{m+1}^{2(r-m)}.$
In view of the integer order ballistic upper bound, for $u\in{\mathcal{H}}^{m+1}({\mathbb{Z}}^{d})$,
(20) $\|e^{-{\rm i}tH}u\|_{r}$ $\leq$ $\|e^{-{\rm i}tH}u\|_{m}^{m+1-r}\|e^{-{\rm i}tH}u\|_{m+1}^{r-m}$
$\leq$ $c^{m+1-r}_{m}c^{r-m}_{m+1}\|u\|_{m}^{m+1-r}\|u\|_{m+1}^{r-m}\cdot t^{\left((m+1)-r\right)m+\left(r-m\right)(m+1)}$
$\leq$ $c_{r}\|u\|_{r}t^{r},\quad{\rm with}\ c_{r}:=c^{m+1-r}_{m}c^{r-m}_{m+1}.$
Note that ${\mathcal{H}}^{m+1}({\mathbb{Z}}^{d})$ is dense in ${\mathcal{H}}^{r}({\mathbb{Z}}^{d})$ in the sense that, given $u\in{\mathcal{H}}^{r}({\mathbb{Z}}^{d})$, for any $\varepsilon>0$, there exists $u_{\varepsilon}\in{\mathcal{H}}^{m+1}({\mathbb{Z}}^{d})$ such that $\|u_{\varepsilon}-u\|_{r}<\varepsilon$. Since (20) implies that, for fixed $t\geq 0$, the time-evolution $e^{-{\rm i}tH}:{\mathcal{H}}^{m+1}({\mathbb{Z}}^{d})\to{\mathcal{H}}^{r}({\mathbb{Z}}^{d})$ is bounded with operator norm controlled by $c_{r}t^{r}$, through the Bounded Linear Transformation Theorem, $e^{-{\rm i}tH}$ is bounded on ${\mathcal{H}}^{r}({\mathbb{Z}}^{d})$ and (16) is satisfied.∎
3.2. Ballistic Lower Bound
Let us next focus on establishing the lower bound in (3).
For the Schrödinger operator $H=-\Delta+V$, with the real potential $V=(V_{n})$ decaying with respect to $|n|$, we show first the following spectral result (compare [8] for a related result under a stronger (short range) decay condition):
Proposition 3.2.
If $V_{n}=o(|n|^{-1})$, then $\sigma_{\rm sc}(H)=\emptyset$ and
$\sigma_{\rm ac}(H)=\sigma(-\Delta)=[-2d,2d].$
Proof.
Since the essential spectrum is invariant under compact perturbations [28], the essential spectrum of $H=-\Delta+V$ is equal to $\sigma_{\rm ess}(-\Delta)=\sigma(-\Delta)=[-2d,2d]$, provided that the potential $V=(V_{n})$ is decaying with respect to $|n|$.
Furthermore, if $V_{n}=o(|n|^{-1})$, then we have the compactness of
$([V,[H,-Q^{2}]]u)_{n}$ $=$ $\sum_{i=1}^{d}\left(\left(2n_{i}+1\right)(V_{n}-V_{n+e_{i}})u_{n+e_{i}}\right$.
$\ \ \ \ \ \ \left.-\left(2n_{i}-1\right)(V_{n}-V_{n-e_{i}})u_{n-e_{i}}\right).$
Hence, given $0<\theta<2d$, for the interval $J_{\theta}$ defined as in (12), we have
$\chi_{J_{\theta}}(H)[H,[H,-Q^{2}]]\chi_{J_{\theta}}(H)\geq\frac{2}{d}\theta\chi_{J_{\theta}}(H)+[V,[H,-Q^{2}]]\chi_{J_{\theta}}(H).$
According to Mourre [26] (see also [2, Section 7.2]), for any $0<\theta<2d$, in $J_{\theta}$ there is no singular continuous spectrum and all eigenvalues are isolated. Since we already know that $\sigma_{\rm ess}(H)=[-2d,2d]$, this forces $\sigma_{\rm ac}(H)=[-2d,2d]$ and concludes the proof of the proposition. ∎
We first show the lower bound in (3) for the case $r=1$. For any non-vanishing $u\in{\mathcal{H}}_{\rm ac}^{1}({\mathbb{Z}}^{d})$, there is $0<\theta<2d$ with $\chi_{J_{\theta}}(H)u$ non-vanishing. According to Corollary 2.9, $e^{-{\rm i}tH}$ exhibits the order1-1 ballistic lower bound: for any non-vanishing $u\in{\mathcal{H}}_{\rm ac}^{1}({\mathbb{Z}}^{d})$, there exists $I_{\theta}\subset J_{\theta}$ with $\mathrm{Leb}(I_{\theta})>0$ and $\chi_{I_{\theta}}(H)u$ non-vanishing, such that there is $a_{1}>0$, independent of $u$, such that
(21) $\|e^{-{\rm i}tH}u\|_{1}>a_{1}\theta^{\frac{1}{2}}\|\chi_{I_{\theta}}(H)u\|_{0}\,t,$
which establishes the lower bound in (3) in the case $r=1$.
The orderr-r ballistic lower bound for $r\geq 1$ is deduced from the following lemma.
Lemma 3.3.
For the self-adjoint operator $H$ on $\ell^{2}({\mathbb{Z}}^{d})$, if there exists $r>0$ such that for any non-vanishing $u\in{\mathcal{H}}_{\rm ac}^{r}({\mathbb{Z}}^{d})$, $\|e^{-itH}u\|_{r}\geq b_{r,u}t^{r}$ for some constant $b_{r,u}>0$, then for any $r^{\prime}>r$ and $u\in{\mathcal{H}}_{\rm ac}^{r^{\prime}}({\mathbb{Z}}^{d})\setminus\{0\}$, there exists a constant $b_{r^{\prime},u}>0$ such that $\|e^{-itH}u\|_{r^{\prime}}\geq b_{r^{\prime},u}t^{r^{\prime}}$.
Remark 3.4.
This argument is a variation of the argument given in the proof of [14, Lemma 2.7].
Proof.
Suppose that $u\in{\mathcal{H}}_{\rm ac}^{r}({\mathbb{Z}}^{d})$ is $\ell^{2}$-normalized. By assumption, there exists a constant $b_{r,u}>0$ such that $\|e^{-itH}u\|_{r}\geq b_{r,u}t^{r}$, which means that
$\sum_{n\in{\mathbb{Z}}^{d}}(1+|n|^{2})^{r}|\langle\delta_{n},e^{-itH}u\rangle|^{2}\geq b^{2}_{r,u}t^{2r}.$
For $r^{\prime}>r$, note that $x\mapsto|x|^{\frac{r^{\prime}}{r}}$ is convex on $\mathbb{R}$}. Using Jensen’s inequality (and emphasizing that $u$ is $\ell^{2}$-normalized so that $n\mapsto|\langle\delta_{n},e^{-itH}u\rangle|^{2}$ is a probability distribution on ${\mathbb{Z}}^{d}$), we observe that, for any $\ell^{2}$-normalized $u\in{\mathcal{H}}_{\rm ac}^{r^{\prime}}({\mathbb{Z}}^{d})\subset{\mathcal{H}}_{\rm ac}^{r}({\mathbb{Z}}^{d})$,
$\sum_{n\in{\mathbb{Z}}^{d}}(1+|n|^{2})^{r^{\prime}}|\langle\delta_{n},e^{-itH}u\rangle|^{2}$ $=\sum_{n\in{\mathbb{Z}}^{d}}\left((1+|n|^{2})^{r}\right)^{\frac{r^{\prime}}{r}}|\langle\delta_{n},e^{-itH}u\rangle|^{2}$
$\geq\left(\sum_{n\in{\mathbb{Z}}^{d}}(1+|n|^{2})^{r}|\langle\delta_{n},e^{-itH}u\rangle|^{2}\right)^{\frac{r^{\prime}}{r}}$
$\geq b^{\frac{2r^{\prime}}{r}}_{r,u}t^{2r^{\prime}},$
which implies that $\|e^{-itH}u\|_{r}\geq b_{r^{\prime},u}t^{r^{\prime}}$ with $b_{r^{\prime},u}=b_{r,u}^{\frac{r^{\prime}}{r}}$. By scaling, the bound extends to non-normalized $u\in{\mathcal{H}}_{\rm ac}^{r^{\prime}}({\mathbb{Z}}^{d})\setminus\{0\}$.∎
As for the case $0<r<1$, we have the following lemma, which completes the proof of Theorem 1.1.
Lemma 3.5.
For the Schrödinger operator $H=-\Delta+V$ with $V_{n}=o(|n|^{-1})$, for any $0<r<1$ and any non-vanishing $u\in{\mathcal{H}}_{\rm ac}^{r}({\mathbb{Z}}^{d})$, there exists a constant $b_{r,u}>0$ such that for $t\geq 1$,
$\|e^{-{\rm i}tH}u\|_{r}\geq b_{r,u}t^{r}.$
Proof.
Similar to (19), applying Hölder’s inequality, we obtain
$\|e^{-{\rm i}tH}u\|_{1}\leq\|e^{-{\rm i}tH}u\|_{r}^{\frac{1}{2-r}}\|e^{-{\rm i}tH}u\|^{\frac{1-r}{2-r}}_{2}.$
According to (17) in Proposition 3.1 and (21), for any $u\in{\mathcal{H}}^{2}_{\rm ac}({\mathbb{Z}}^{d})\setminus\{0\}$, there exist $0<\theta<2d$ and $I_{\theta}\subset J_{\theta}$, with $J_{\theta}$ defined as in (12), such that $\chi_{I_{\theta}}(H)u$ is non-vanishing and
$\|e^{-{\rm i}tH}u\|_{1}\geq a_{1}\theta^{\frac{1}{2}}\|\chi_{I_{\theta}}(H)u\|_{0}\,t,\quad\|e^{-{\rm i}tH}u\|_{2}\leq\|u\|_{2}+\|u\|_{1}t+\tilde{c}_{2}\|u\|_{0}\,t^{2},$
which implies that, as $t\to\infty$,
$\|e^{-{\rm i}tH}u\|_{r}\geq\frac{(a_{1}\theta^{\frac{1}{2}}\|\chi_{I_{\theta}}(H)u\|_{0}t)^{2-r}}{(\|u\|_{2}+\|u\|_{1}t+\tilde{c}_{2}\|u\|_{0}t^{2})^{1-r}}\geq\frac{(a_{1}\theta^{\frac{1}{2}}\|\chi_{I_{\theta}}(H)u\|_{0})^{2-r}}{2^{1-r}(\tilde{c}_{2}\|u\|_{0})^{1-r}}t^{r}.$
Note that ${\mathcal{H}}^{2}_{\rm ac}({\mathbb{Z}}^{d})={\mathcal{H}}^{2}({\mathbb{Z}}^{d})\cap{\mathcal{H}}_{\rm ac}(H)$ is dense in ${\mathcal{H}}^{r}_{\rm ac}({\mathbb{Z}}^{d})={\mathcal{H}}^{r}({\mathbb{Z}}^{d})\cap{\mathcal{H}}_{\rm ac}(H)$ in the sense that, given $u\in{\mathcal{H}}_{\rm ac}^{r}({\mathbb{Z}}^{d})$, for any $\varepsilon>0$, there exists $u_{\varepsilon}\in{\mathcal{H}}_{\rm ac}^{2}({\mathbb{Z}}^{d})$ such that $\|u_{\varepsilon}-u\|_{r}<\varepsilon$, which implies that, for $\varepsilon$ sufficiently small,
$\|\chi_{I_{\theta}}(H)u_{\varepsilon}\|_{0}$ $\geq$ $\|\chi_{I_{\theta}}(H)u\|_{0}-\|u_{\varepsilon}-u\|_{0}$
$\geq$ $\|\chi_{I_{\theta}}(H)u\|_{0}-\|u_{\varepsilon}-u\|_{r}\ \geq\ \frac{1}{2}\|\chi_{I_{\theta}}(H)u\|_{0},$
$\|u_{\varepsilon}\|_{0}$ $\leq$ $\|u\|_{0}+\|u_{\varepsilon}-u\|_{0}\ \leq\ 2\|u\|_{0}.$
According to (16) in Proposition 3.1, for $t\geq 0$,
$\|e^{-{\rm i}tH}(u_{\varepsilon}-u)\|_{r}\leq c_{r}\|u_{\varepsilon}-u\|_{r}t^{r}.$
Then, for $\varepsilon$ sufficiently small, as $t\to\infty$, we have
$\|e^{-{\rm i}tH}u\|_{r}$ $\geq\|e^{-{\rm i}tH}u_{\varepsilon}\|_{r}-\|e^{-{\rm i}tH}(u_{\varepsilon}-u)\|_{r}$
$\geq\frac{(a_{1}\theta^{\frac{1}{2}}\|\chi_{I_{\theta}}(H)u_{\varepsilon}\|_{0})^{2-r}}{2^{1-r}(\tilde{c}_{2}\|u_{\varepsilon}\|_{0})^{1-r}}\cdot t^{r}-c_{r}\|u_{\varepsilon}-u\|_{r}t^{r}$
$\geq\left(\frac{(a_{1}\theta^{\frac{1}{2}}\|\chi_{I_{\theta}}(H)u\|_{0})^{2-r}}{2^{4-3r}(\tilde{c}_{2}\|u\|_{0})^{1-r}}-c_{r}\varepsilon\right)t^{r}$
$\geq\frac{(a_{1}\theta^{\frac{1}{2}}\|\chi_{I_{\theta}}(H)u\|_{0})^{2-r}}{16(\tilde{c}_{2}\|u\|_{0})^{1-r}}t^{r}.\qed$
Acknowledgment
We are grateful to Sergey Denisov, Jake Fillman, and Tal Malinovitch for helpful information about the literature.