### v3

# Ballistic Transport for Discrete Multi-Dimensional Schrödinger Operators With Decaying Potential

Abstract.
We consider the discrete Schrödinger operator $H=-\Delta+V$ on $\ell^{2}({\mathbb{Z}}^{d})$ with a decaying potential, in arbitrary lattice dimension $d\in{\mathbb{N}}^{*}$, where $\Delta$ is the standard discrete Laplacian and $V_{n}=o(|n|^{-1})$ as $|n|\to\infty$. We prove the absence of singular continuous spectrum for $H$. For the unitary evolution $e^{-{\rm i}tH}$, we prove that it exhibits ballistic transport in the sense that, for any $r>0$, the weighted $\ell^{2}$-norm
$\|e^{-{\rm i}tH}u\|_{r}:=\left(\sum_{n\in{\mathbb{Z}}^{d}}(1+|n|^{2})^{r}|(e^{-{\rm i}tH}u)_{n}|^{2}\right)^{\frac{1}{2}}$
grows at rate $\simeq t^{r}$ as $t\to\infty$, provided that the initial state $u$ is in the absolutely continuous subspace and satisfies $\|u\|_{r}<\infty$.
The proof relies on commutator methods and a refined Mourre estimate, which yields quantitative lower bounds on transport for operators with purely absolutely continuous spectrum over appropriate spectral intervals. Compactness arguments and localized spectral projections are used to extend the result to perturbed operators, extending the classical result for the free Laplacian to a broader class of decaying potentials.
Key words and phrases:
Schrödinger operators, ballistic transport, Mourre estimate
D. D. was supported in part by NSF grants DMS–2054752 and DMS–2349919
Z. Zhao was partially supported by ANR grant for ANR-22-CE40-0016 and partially supported by NSFC grant 12471178.
1. Introduction and Main Result
1.1. General Setting and Questions
A main motivation to perform a spectral analysis of a Schrödinger operator $H=-\Delta+V$ is given by the known connections between spectral properties of $H$ and the long-term asymptotics of the associated unitary group $e^{-{\rm i}tH}$, with which one can describe the solutions of the associated time-dependent Schrödinger equation. That is, $\psi(t)=e^{-{\rm i}tH}\psi_{0}$ is the unique solution of the time-dependent equation
${\rm i}\partial_{t}\psi=H\psi,\qquad\left.\psi\right|_{t=0}=\psi_{0}.$
A fundamental result establishing such a connection is given by the RAGE theorem; compare, for example, [10, Theorem 1.6.7]. More refined connections can be found in [10, 25] and references therein.
Generally speaking, the more regular the spectral measure $\mu=\mu_{H,\psi_{0}}$ associated with the pair $(H,\psi_{0})$ is, the greater the tendency of $\psi$ is to leave compact regions in space as time grows.
In order to be more specific, let us consider the case where the operator $H$ acts in the Hilbert space $\ell^{2}({\mathbb{Z}}^{d})$; analogous statements exist in the case of the continuum, $L^{2}({\mathbb{R}}^{d})$, or the half line, $\ell^{2}({\mathbb{Z}}_{+})$ or $L^{2}({\mathbb{R}}_{+})$.
Consider the Lebesgue decomposition of $\mu$ into its absolutely continuous, singular continuous, and pure point parts:
$\mu=\mu_{\mathrm{ac}}+\mu_{\mathrm{sc}}+\mu_{\mathrm{pp}}.$
Let us first discuss the question of whether the time evolution leaves compact regions or is confined to a suitable compact region, up to an arbitrarily small error. The RAGE theorem states that
$\mu=\mu_{\mathrm{pp}}$ $\Longleftrightarrow$ $\forall\ \varepsilon>0,\ \exists\ N,\ \forall\ t\in{\mathbb{R}}\;:\;\sum_{|n|>N}|\langle\delta_{n},\psi(t)\rangle|^{2}<\varepsilon,$
$\mu=\mu_{\mathrm{sc}}+\mu_{\mathrm{ac}}$ $\Longleftrightarrow$ $\forall\ N\;:\;\lim_{T\to\infty}\frac{1}{2T}\int_{-T}^{T}\sum_{|n|\leq N}|\langle\delta_{n},\psi(t)\rangle|^{2}\,dt=0,$
$\mu=\mu_{\mathrm{ac}}$ $\Longrightarrow$ $\forall\ N\;:\;\lim_{|t|\to\infty}\sum_{|n|\leq N}|\langle\delta_{n},\psi(t)\rangle|^{2}=0.$
Recall that for a normalized initial state, $\|\psi_{0}\|=1$, $\sum_{|n|\leq N}|\langle\delta_{n},\psi(t)\rangle|^{2}$ gives the probability of the wavefunction being in the ball of radius $N$ centered at the origin at time $t$, so these statements indeed have the interpretation alluded to above.
We will be especially interested in cases where the spectral measure is absolutely continuous (a.c. for short). This is in response to the following general question: for initial states $\psi_{0}$ with $\mu=\mu_{\mathrm{ac}}$, is transport ballistic (formulated in Subsection 1.3)? There are general remarks that are relevant to this question (see the survey [13] for a much more in-depth discussion):
  • (i)
    One expects this to be true for many models of interest, but there is no general result.
  • (ii)
    In fact, there cannot be a completely general result as there are counterexamples, which can be constructed via inverse spectral theory in higher dimensions; see Bellissard and Schulz-Baldes [4].
Whether or not $\mu=\mu_{\mathrm{ac}}$ implies ballistic transport in one dimension being a difficult question in general, it is natural to explore it for specific interesting classes of potentials. This has indeed been done. For periodic $V$, it was shown by Asch and Knauf [3] in the continuum setting and Damanik, Lukic, and Yessen [12] in the discrete case. An extension to suitable limit-periodic potentials was obtained by Fillman [16] in the discrete case and by Young [36] in the continuum. Finally, suitable quasi-periodic potentials were studied by Zhao [38, 39], Zhang and Zhao [37], and Ge and Kachkovskiy [19]. In higher dimensions, there are results of Asch and Knauf [3], Black, Damanik, Malinovitch, and Young [6], Boutet de Monvel and Sabri [7], Fillman [17], and Karpeshina, Lee, Parnovski, Shterenberg, and Stolz [23, 24]. [^1] It should be noted, however, that [23, 24] consider the time-averaged evolution, whereas this paper and the other papers mentioned above do not.
While periodic, limit-periodic, and quasi-periodic potentials are of course of great importance with respect to actual physical models, another very important class is that of decaying potentials, and the absence of a ballistic transport result for them is quite striking. The present paper seeks to address this and produce a ballistic transport result for suitable decaying potentials.
1.2. Absence of Singular Continuous Spectrum
Consider the $d$-dimensional discrete Schrödinger operator $H_{V}=-\Delta+V$ defined through the discrete Laplacian
(1) $\Delta:\ell^{2}({\mathbb{Z}}^{d})\to\ell^{2}({\mathbb{Z}}^{d}),\quad(\Delta u)_{n}=\sum_{j=1}^{d}(\Delta_{j}u)_{n}:=\sum_{j=1}^{d}(u_{n+e_{j}}+u_{n-e_{j}}),$
with $\{e_{j}\}_{j=1}^{d}$ the canonical basis of ${\mathbb{Z}}^{d}$, and the multiplication by the potential $V=(V_{n})$.
Our first main result is about the spectral type of $H_{V}$.
Theorem 1.1.
If $V=(V_{n})$ satisfies
(2) $V_{n}=o(|n|^{-1}),\quad|n|\to\infty,$
then the Schrödinger operator $H_{V}=-\Delta+V$ has no singular continuous spectrum. Moreover, for any closed interval ${\mathcal{I}}\subset(-2d,2d)$, there are at most finitely many eigenvalues of $H_{V}$ in $\mathcal{I}$} and each of these eigenvalues is of finite multiplicity.
The decay condition (2) of $V$ arises naturally from our proof. For a conjugate operator $A$ which establishes the Mourre estimate (see Section 2) for $H_{V}$, the condition (2) implies that $H_{V}\in C^{1}(A)$. Under a stronger short-range decay condition of $V$, the absence of singular continuous (s.c. for short) spectrum for $H_{V}$ is demonstrated in [8]. This potential ensures that $H_{V}$ belongs to the class $\mathscr{C}^{1,1}(A)\subset C^{1}(A)$ (as defined in [2, Definition 6.2.2]). Additionally, it is noted in [8, 3.3] that it remains an open question whether the $C^{1}(A)$ regularity property is sufficient for the absence of s.c. spectrum of $H_{V}$ in the subset on which the Mourre estimate holds. See Section 3 for more detailed discussions.
In addition, the condition (2) is known to be the sharp decay condition that ensures that the interior of the essential spectrum is purely a.c. in one dimension [35]. Note that Wigner-von Neumann potentials satisfy $V_{n}=O(|n|^{-1})$, $|n|\to\infty$, and allow for eigenvalues embedded in the interior of the essential spectrum [11, 34]. [^2] See especially [11, Example on p.52] for how to embed an eigenvalue in the discrete case. Whether or not (2) ensures purely a.c. interior essential spectrum in higher dimensions as well is not known (to the best of our knowledge). At this point we cannot rule out the existence of isolated eigenvalues embedded in the interior of the essential spectrum.
1.3. Ballistic Transport
For $r\in{\mathbb{R}}_{+}$ and $d\in{\mathbb{N}}^{*}$, let us define the subspace ${\mathcal{H}}^{r}({\mathbb{Z}}^{d})$ of $\ell^{2}({\mathbb{Z}}^{d})$, as
${\mathcal{H}}^{r}({\mathbb{Z}}^{d}):=\left\{u=(u_{n})\in\ell^{2}({\mathbb{Z}}^{d}):\sum_{n\in{\mathbb{Z}}^{d}}(1+|n|^{2})^{r}|u_{n}|^{2}<\infty\right\},$
with $|n|:=\sqrt{n_{1}^{2}+\cdots+n_{d}^{2}}$. It also naturally defines the weighted $\ell^{2}$-norm $\|\cdot\|_{r}$ on ${\mathcal{H}}^{r}({\mathbb{Z}}^{d})$ via
$\|u\|_{r}:=\left(\sum_{n\in{\mathbb{Z}}^{d}}(1+|n|^{2})^{r}|u_{n}|^{2}\right)^{1/2}.$
We identify the $\|\cdot\|_{\ell^{2}}$-norm with the $\|\cdot\|_{0}$-norm.
Let ${\mathcal{H}}_{\rm ac}(H_{V})$ be the absolutely continuous spectral subspace associated with $H_{V}$. For $r\geq 0$, let ${\mathcal{H}}_{\rm ac}^{r}({\mathbb{Z}}^{d}):={\mathcal{H}}^{r}({\mathbb{Z}}^{d})\cap{\mathcal{H}}_{\rm ac}(H_{V})$. Here is our main result:
Theorem 1.2.
If $V=(V_{n})$ satisfies (2), then the unitary evolution $e^{-{\rm i}tH_{V}}$ exhibits ballistic transport in the sense that, for any $u\in{\mathcal{H}}_{\rm ac}^{r}({\mathbb{Z}}^{d})\setminus\{0\}$ and any $r>0$, there exists a constant $C_{u,r}>1$ such that
(3) $C_{u,r}^{-1}\leq\frac{\|e^{-{\rm i}tH_{V}}u\|_{r}}{t^{r}}\leq C_{u,r}\quad\forall\ t\geq 1.$
Remark 1.3.
In the terminology of the survey [13], Theorem 1.2 establishes ballistic transport in the norm sense for all moments under the decay assumption (2).
Clearly, the ballistic lower bound in (3) cannot hold for eigenstates, and hence we naturally restrict our attention in Theorem 1.2 to initial states in the absolutely continuous subspace.
The remainder of the paper is organized as follows. As a key ingredient of the proof, the Mourre estimate for $H_{V}$ is established in Section 2. Based on this estimate, Theorem 1.1 is proved in Section 3, and the order1-1 ballistic lower bound is derived in Section 4. Finally, the proof of Theorem 1.2 is completed in Section 5, where ballistic upper and lower bounds of arbitrary order are established.
2. Mourre estimate
Define the weight operator $Q:{\mathcal{H}}^{1}({\mathbb{Z}}^{d})\to\ell^{2}({\mathbb{Z}}^{d})$ by
(4) $(Qu)_{n}:=\sqrt{1+|n|^{2}}\,u_{n},\quad n\in{\mathbb{Z}}^{d}.$
We introduce the conjugate operator
$A:=-{\rm i}[H_{V},-Q^{2}]={\rm i}[Q^{2},\Delta],$
which for $u\in D(A)$ can be written explicitly as
$(Au)_{n}=-{\rm i}\sum_{j=1}^{d}\left((2n_{j}+1)u_{n+e_{j}}-(2n_{j}-1)u_{n-e_{j}}\right),\quad n\in{\mathbb{Z}}^{d}.$
It is straightforward to verify that ${\mathcal{H}}^{1}({\mathbb{Z}}^{d})\subset D(A)$ and that
$\|Au\|_{0}\leq 6d\,\|u\|_{1},\quad u\in{\mathcal{H}}^{1}({\mathbb{Z}}^{d}).$
In particular, $D(A)$ is dense in $\ell^{2}({\mathbb{Z}}^{d})$.
Since the essential spectrum is invariant under compact perturbations (see [33]), the essential spectrum of $H_{V}$ coincides with that of $H_{0}:=-\Delta$, namely
$\sigma_{\rm ess}(H_{V})=\sigma(H_{0})=[-2d,2d].$
Indeed, under assumption (2), the multiplication operator $V=(V_{n})$ is compact on $\ell^{2}({\mathbb{Z}}^{d})$, since $V_{n}\to 0$ as $|n|\to\infty$.
A further computation shows that, for $u\in D(A)$,
(5) ${\rm i}[H_{0},A]u$ $=$ $2\sum_{j=1}^{d}(4{\rm I}-\Delta_{j}^{2})u,$
$({\rm i}[V,A]u)_{n}$ $=$ $\sum_{j=1}^{d}\left(\left(2n_{j}+1\right)(V_{n}-V_{n+e_{j}})u_{n+e_{j}}\right$.
$\ \ \ \ \ \ \left.-\left(2n_{j}-1\right)(V_{n}-V_{n-e_{j}})u_{n-e_{j}}\right).$
Under assumption (2), the operator ${\rm i}[V,A]$ extends to a compact operator on $\ell^{2}({\mathbb{Z}}^{d})$. Consequently, ${\rm i}[H_{V},A]$ extends to a bounded operator on $\ell^{2}({\mathbb{Z}}^{d})$. In the framework of Amrein, Boutet de Monvel and Georgescu, this implies that $H_{V}\in C^{1}(A)$ (see [2, Definition 6.2.2 and Theorem 6.2.10] and [8, 3.2.2] for precise definitions and properties).
Define the set of thresholds
${\mathcal{Z}}:=[-2d,2d]\cap\{2d+4k\}_{k\in{\mathbb{Z}}}.$
Then the set $2d,2d]\setminus{\mathcal{Z}}$ consists of $d$ disjoint open intervals, each of length $4$. We denote these intervals by ${\mathcal{J}}_{j}$, $1\leq j\leq d$.
Lemma 2.1.
Let $E\in[-2d,2d]\setminus{\mathcal{Z}}$. Then there exists $\delta_{0}>0$ such that, for ${\mathcal{I}}_{0}=(E-\delta_{0},E+\delta_{0})$, there exists $0<\alpha<8d^{-2}$ satisfying
(7) $\chi_{{\mathcal{I}}_{0}}(H_{V}){\rm i}[H_{V},A]\chi_{{\mathcal{I}}_{0}}(H_{V})\geq\alpha\,\chi_{{\mathcal{I}}_{0}}(H_{V})+\chi_{{\mathcal{I}}_{0}}(H_{V})K_{V}\chi_{{\mathcal{I}}_{0}}(H_{V}),$
where $K_{V}$ is a compact operator on $\ell^{2}({\mathbb{Z}}^{d})$.
Proof.
Let $E\in{\mathcal{J}}_{j}\subset[-2d,2d]\setminus{\mathcal{Z}}$ for some $1\leq j\leq d$. Then a sufficiently small neighborhood of $E$ is still contained in ${\mathcal{J}}_{j}$. Moreover, there exists $\delta_{0}>0$ such that
(8) ${\rm dist}((E-2\delta_{0},E+2\delta_{0}),{\mathcal{Z}})>\frac{d\sqrt{\alpha}}{\sqrt{2}}\ {\rm for\ some}\ 0<\alpha<8d^{-2}.$
Under the Fourier transform
$\widehat{u}(\theta)=\sum_{n\in{\mathbb{Z}}^{d}}u_{n}e^{{\rm i}\langle n,\theta\rangle},\quad\theta\in{\mathbb{T}}^{d}=[0,2\pi)^{d},$
the operator $H_{0}=-\Delta$ becomes a multiplication operator:
$\widehat{(-\Delta u)}(\theta)=-\left(\sum_{j=1}^{d}\left(e^{{\rm i}\theta_{j}}+e^{-{\rm i}\theta_{j}}\right)\right)\widehat{u}(\theta)=-\left(2\sum_{j=1}^{d}\cos(\theta_{j})\right)\widehat{u}(\theta).$
The thresholds in $\mathcal{Z}$} correspond to the values
$E(\theta):=-2\sum_{j=1}^{d}\cos(\theta_{j})\quad{\rm with}\ \theta_{j}\in\{0,\pi\}.$
For $\theta$ such that $E(\theta)\in{\mathcal{I}}_{0}^{\prime}:=(E-2\delta_{0},E+2\delta_{0})$ and (8) holds, there exists $1\leq j_{*}\leq d$ such that
$|\cos(\theta_{j_{*}})+1|,\ |\cos(\theta_{j_{*}})-1|>\frac{\sqrt{\alpha}}{2\sqrt{2}}.$
In view of (5), we have
$\widehat{({\rm i}[H_{0},A]u)}(\theta)=\left(8\sum_{j=1}^{d}\sin^{2}(\theta_{j})\right)\widehat{u}(\theta),$
which implies
$\chi_{{\mathcal{I}}_{0}^{\prime}}(H_{0}){\rm i}[H_{0},A]\chi_{{\mathcal{I}}_{0}^{\prime}}(H_{0})\geq 8\left(\frac{\sqrt{\alpha}}{2\sqrt{2}}\right)^{2}\chi_{{\mathcal{I}}_{0}^{\prime}}(H_{0})=\alpha\chi_{{\mathcal{I}}_{0}^{\prime}}(H_{0}).$
Let $g_{{\mathcal{I}}_{0}^{\prime}}\in C_{c}^{\infty}({\mathbb{R}},[0,1])$ be supported on ${\mathcal{I}}_{0}^{\prime}$ and identically $1$ on ${\mathcal{I}}_{0}=(E-\delta_{0},E+\delta_{0})$. Noting that
$g_{{\mathcal{I}}_{0}^{\prime}}(H_{0})=g_{{\mathcal{I}}_{0}^{\prime}}(H_{0})\chi_{{\mathcal{I}}_{0}^{\prime}}(H_{0})=\chi_{{\mathcal{I}}_{0}^{\prime}}(H_{0})g_{{\mathcal{I}}_{0}^{\prime}}(H_{0}),$
and therefore
$g_{{\mathcal{I}}_{0}^{\prime}}(H_{0}){\rm i}[H_{0},A]g_{{\mathcal{I}}_{0}^{\prime}}(H_{0})\geq\alpha g^{2}_{{\mathcal{I}}_{0}^{\prime}}(H_{0}).$
Since $V$ is a multiplication operator with $V_{n}\to 0$, it is compact on $\ell^{2}({\mathbb{Z}}^{d})$. Hence $(H_{V}-z)^{-1}-(H_{0}-z)^{-1}$ is compact for $z\notin{\mathbb{R}}$. By the Helffer–Sjöstrand functional calculus [22], it follows that $g_{{\mathcal{I}}_{0}^{\prime}}(H_{V})-g_{{\mathcal{I}}_{0}^{\prime}}(H_{0})$ and $g^{2}_{{\mathcal{I}}_{0}^{\prime}}(H_{V})-g^{2}_{{\mathcal{I}}_{0}^{\prime}}(H_{0})$ are compact. Moreover, under assumption (2), the commutator ${\rm i}[V,A]$ is compact. It follows that
$g_{{\mathcal{I}}_{0}^{\prime}}(H_{V}){\rm i}[H_{V},A]g_{{\mathcal{I}}_{0}^{\prime}}(H_{V})\geq\alpha g^{2}_{{\mathcal{I}}_{0}^{\prime}}(H_{V})+K_{V},$
where $K_{V}$ is the compact operator defined by
$K_{V}$ $:=$ $g_{{\mathcal{I}}_{0}^{\prime}}(H_{V}){\rm i}[V,A]g_{{\mathcal{I}}_{0}^{\prime}}(H_{V})$
$+\,(g_{{\mathcal{I}}_{0}^{\prime}}(H_{V})-g_{{\mathcal{I}}_{0}^{\prime}}(H_{0})){\rm i}[H_{0},A]g_{{\mathcal{I}}_{0}^{\prime}}(H_{V})$
$+\,g_{{\mathcal{I}}_{0}^{\prime}}(H_{0}){\rm i}[H_{0},A](g_{{\mathcal{I}}_{0}^{\prime}}(H_{V})-g_{{\mathcal{I}}_{0}^{\prime}}(H_{0}))$
$-\,\alpha(g^{2}_{{\mathcal{I}}_{0}^{\prime}}(H_{V})-g^{2}_{{\mathcal{I}}_{0}^{\prime}}(H_{0})).$
Each term in $K_{V}$ is compact, being a product of bounded and compact operators. Since $g_{{\mathcal{I}}_{0}^{\prime}}\equiv 1$ on ${\mathcal{I}}_{0}=(E-\delta_{0},E+\delta_{0})$, we have
$\chi_{{\mathcal{I}}_{0}}(H_{V}){\rm i}[H_{V},A]\chi_{{\mathcal{I}}_{0}}(H_{V})$
$=$ $\chi_{{\mathcal{I}}_{0}}(H_{V})g_{{\mathcal{I}}_{0}^{\prime}}(H_{V}){\rm i}[H_{V},A]g_{{\mathcal{I}}_{0}^{\prime}}(H_{V})\chi_{{\mathcal{I}}_{0}}(H_{V})$
$>$ $\alpha\chi_{{\mathcal{I}}_{0}}(H_{V})+\chi_{{\mathcal{I}}_{0}}(H_{V})K_{V}\chi_{{\mathcal{I}}_{0}}(H_{V}),$
which proves (7). ∎
Remark 2.2.
In view of the above proof, we observe that the parameter $\delta_{0}$ in (7) depends only on the distance of $E$ to the threshold set $\mathcal{Z}$} (and the choice of $\alpha$), and is independent of the potential $V$.
3. Absence of Singular Continuous Spectrum
It is well known that the Mourre estimate (7) in Lemma 2.1 implies the absence of singular continuous (s.c.) spectrum, under suitable regularity assumptions on $H_{V}$ with respect to $A$.
Theorem 3.1.
For $E\in[-2d,2d]\setminus{\mathcal{Z}}$, there is $\delta>0$ such that
  1. (i)
    $(E-\delta,E+\delta)\subset[-2d,2d]\setminus{\mathcal{Z}}$, and $H_{V}$ has at most finitely many eigenvalues in $(E-\delta,E+\delta)$, each of finite multiplicity.
  2. (ii)
    For each closed interval
    $a,b]\subset(E-\delta,E+\delta)\,\cap\,[-2d,2d]\setminus({\mathcal{Z}}\cup\sigma_{\rm pp}(H_{V})),$
    there exists a constant $c_{0}>0$ such that, for $\langle A\rangle:=(A^{2}+{\rm I})^{\frac{1}{2}}$,
    (9) $\sup_{\Re z\in[a,b]\atop{\Im z\neq 0}}\|\langle A\rangle^{-1}(H_{V}-z)^{-1}\langle A\rangle^{-1}\|_{0}<c_{0},$
    where $\|\cdot\|_{0}$ denotes the operator norm on $\ell^{2}({\mathbb{Z}}^{d})$.
The estimate (9) is commonly referred to as the limiting absorption principle (LAP) for the operator $H_{V}$. As a consequence [2, Proposition 7.1.3], $\sigma_{\rm sc}(H_{V})=\varnothing$ and $\sigma_{\rm ac}(H_{V})=[-2d,2d]$.
Remark 3.2.
Beyond the setting of $\ell^{2}({\mathbb{Z}}^{d})$, Mourre estimates have also been employed to establish the absence of s.c. spectrum for adjacency operators on graphs; see, for instance, [27]. Moreover, Mourre theory plays a central role in scattering theory for Schrödinger operators as well as for a broader class of dispersive PDEs; see, e.g., [5, 15, 18, 20, 28, 29, 31].
Since $H_{V}\in C^{1}(A)$, assertion (i) of Theorem 3.1 follows from [2, Corollary 7.2.11]. The remainder of this section is devoted to the proof of assertion (ii) of Theorem 3.1.
The absence of s.c. spectrum was established in [8] under a stronger short-range decay assumption on the potential $V$, which ensures that $H_{V}$ belongs to the class $\mathscr{C}^{1,1}(A)\subset C^{1}(A)$ (see [2, Definition 6.2.2]). According to [2, Theorem 7.4.2], if $H_{V}\in\mathscr{C}^{1,1}(A)$ and the Mourre estimate (7) holds, then $H_{V}$ has no s.c. spectrum. It is further pointed out in [8, Section 3.3] that it remains an open question whether the weaker regularity assumption $C^{1}(A)$ suffices to ensure the absence of singular continuous spectrum for $H_{V}$ on the set where (7) holds.
It is instructive to contrast this with the framework developed in [21, 26], which addresses discrete Schrödinger operators perturbed by Wigner–von Neumann type and long-range oscillatory potentials. The analysis is based on a weighted Mourre estimate, in which the commutator is modified by suitable functions of the conjugate operator. This approach allows one to establish LAP under significantly weaker decay and regularity assumptions.
3.1. Quantitative LAP for Rapidly Decaying Potential
Consider the operator $H_{W}$ with a rapidly decaying potential $W$ satisfying
$W_{n}=O(|n|^{-2}).$
In the framework of [2], one has $H_{W}\in C^{2}(A)\subset\mathscr{C}^{1,1}(A)$. In particular, $H_{W}$ has no s.c. spectrum. We shall establish a quantitative LAP following the original approach of Mourre [30].
Let $E\in[-2d,2d]\setminus({\mathcal{Z}}\cup\sigma_{\rm pp}(H_{W}))$, and let ${\mathcal{I}}_{0}=(E-\delta_{0},E+\delta_{0})$ be an open interval on which the Mourre estimate holds:
$\chi_{{\mathcal{I}}_{0}}(H_{W})\,{\rm i}[H_{W},A]\,\chi_{{\mathcal{I}}_{0}}(H_{W})\geq\alpha\chi_{{\mathcal{I}}_{0}}(H_{W})+\chi_{{\mathcal{I}}_{0}}(H_{W})\,K_{W}\,\chi_{{\mathcal{I}}_{0}}(H_{W}),$
with $\alpha>0$ and $K_{W}$ compact, as in Lemma 2.1. Since $E\not\in\sigma_{\rm pp}(H_{W})$, the spectral projections $\chi_{(E-\delta,E+\delta)}(H_{W})$ converge weakly to 0 as $\delta\to 0$. Consequently, one can choose
  • $0<\delta_{1}<\frac{\delta_{0}}{2}$ and define the intervals
    ${\mathcal{I}}^{\prime}_{1}=(E-2\delta_{1},E+2\delta_{1}),\qquad{\mathcal{I}}_{1}=(E-\delta_{1},E+\delta_{1}),$
  • a function $g\in C_{c}^{\infty}({\mathbb{R}},[0,1])$ supported in ${\mathcal{I}}^{\prime}_{1}$ and identically equal to $1$ on ${\mathcal{I}}_{1}$,
such that
$\pm\,g(H_{W})K_{W}g(H_{W})\leq\frac{\alpha}{2}g^{2}(H_{W}).$
It follows that
(10) ${\mathcal{B}}_{W}:=g(H_{W}){\rm i}[H_{W},A]g(H_{W})\geq\frac{\alpha}{2}g^{2}(H_{W}).$
Let us define the quantities
(11) $m_{W}:=\sup_{n\in{\mathbb{Z}}^{d}}|n||W_{n}|,\qquad M_{W}:=\sup_{n\in{\mathbb{Z}}^{d}}|n|^{2}|W_{n}|.$
In view of (5) and (2), there exists a constant $c_{1}>0$ such that
(12) $\|{\mathcal{B}}_{W}\|_{0}\leq\|g(H_{W})\|_{0}^{2}\left(\|[H_{0},A]\|_{0}+\|[W,A]\|_{0}\right)\leq c_{1}(1+m_{W}).$
By [30, Proposition II.5], the operator $G^{W}_{z}(\varepsilon):=(H_{W}-z-{\rm i}\varepsilon{\mathcal{B}}_{W})^{-1}$ is well defined whenever $\Im z\cdot\varepsilon>0$. In the following, we assume that $\Im z>0$ and $\varepsilon>0$. Then there exists a constant $c_{2}>0$, depending only on $\alpha$, such that
(13) $\|G^{W}_{z}(\varepsilon)\|_{0}\leq c_{2}\,\varepsilon^{-1},\quad 0<\varepsilon<1.$
Throughout this section, with $\alpha>0$ fixed, all implicit constants appearing in estimates denoted by $\lesssim$” and $\gtrsim$” are strictly positive and uniform in $\varepsilon$, $z$, and $W$ within $W_{n}=O(|n|^{-2})$.
Consider the auxiliary operator
$F^{W}_{z}(\varepsilon):=\langle A\rangle^{-1}G^{W}_{z}(\varepsilon)\langle A\rangle^{-1}.$
According to [30, Proposition II.5], one has
$\|g(H_{W})G^{W}_{z}(\varepsilon)\langle A\rangle^{-1}\|_{0}\lesssim\|F^{W}_{z}(\varepsilon)\|_{0}^{\frac{1}{2}}\,\varepsilon^{-\frac{1}{2}}.$
Moreover, using the resolvent identity
$G^{W}_{z}(\varepsilon)=G^{W}_{z}(0)\bigl(1+{\rm i}\varepsilon{\mathcal{B}}_{W}G^{W}_{z}(\varepsilon)\bigr),$
we have, in view of (12) and (13),
$\|(1-g(H_{W}))G^{W}_{z}(\varepsilon)\langle A\rangle^{-1}\|_{0}$
$\leq$ $\|(1-g(H_{W}))G^{W}_{z}(0)\|_{0}\cdot\|(1+{\rm i}\varepsilon{\mathcal{B}}_{W}G^{W}_{z}(\varepsilon))\langle A\rangle^{-1}\|_{0}$
$\leq$ $(1+\varepsilon\|{\mathcal{B}}_{W}\|_{0}\|G^{W}_{z}(\varepsilon)\|_{0})\cdot\|(1-g(H_{W}))G^{W}_{z}(0)\|_{0}$
$\lesssim$ $(1+m_{W})\|(1-g(H_{W}))G^{W}_{z}(0)\|_{0}.$
Restricting $\Re z$ to a compact interval $a,b]\subset(E-\delta_{1},E+\delta_{1})$, the operator $(1-g(H_{W}))G^{W}_{z}(0)$ is uniformly bounded (with a bound depending only on $\delta_{1}$). Consequently,
(14) $\|G^{W}_{z}(\varepsilon)\langle A\rangle^{-1}\|_{0}\lesssim 1+m_{W}+\|F^{W}_{z}(\varepsilon)\|_{0}^{\frac{1}{2}}\,\varepsilon^{-\frac{1}{2}}.$
Similarly,
(15) $\|\langle A\rangle^{-1}G^{W}_{z}(\varepsilon)\|_{0}\lesssim 1+m_{W}+\|F^{W}_{z}(\varepsilon)\|_{0}^{\frac{1}{2}}\,\varepsilon^{-\frac{1}{2}}.$
Assuming that $m_{W}\lesssim 1$, we have the a priori bound
(16) $\|F^{W}_{z}(\varepsilon)\|_{0}\lesssim\varepsilon^{-1},\qquad\Re z\in[a,b].$
In view of [30, P. 405], one has
$\frac{d}{d\varepsilon}F^{W}_{z}(\varepsilon)=\langle A\rangle^{-1}G^{W}_{z}(\varepsilon)g(H_{W}){\rm i}[H_{W},A]g(H_{W})G^{W}_{z}(\varepsilon)\langle A\rangle^{-1}.$
Consequently,
(17) $\left\|\frac{d}{d\varepsilon}F^{W}_{z}(\varepsilon)\right\|_{0}$ $\lesssim$ $1+\|F^{W}_{z}(\varepsilon)\|^{\frac{1}{2}}_{0}\varepsilon^{-\frac{1}{2}}$
$+\,\left\|\langle A\rangle^{-1}G^{W}_{z}(\varepsilon){\rm i}[H_{W},A]G^{W}_{z}(\varepsilon)\langle A\rangle^{-1}\right\|_{0}.$
For every $z$ with $\Im z>0$ and $\varepsilon>0$, we write
(19) $\langle A\rangle^{-1}G^{W}_{z}(\varepsilon){\rm i}[H_{W},A]G^{W}_{z}(\varepsilon)\langle A\rangle^{-1}$
$=$ $\langle A\rangle^{-1}G^{W}_{z}(\varepsilon){\rm i}\left[H_{W}-z-{\rm i}\varepsilon{\mathcal{B}}_{W},A\right]G^{W}_{z}(\varepsilon)\langle A\rangle^{-1}$
$-\,\varepsilon\langle A\rangle^{-1}G^{W}_{z}(\varepsilon)\left[{\mathcal{B}}_{W},A\right]G^{W}_{z}(\varepsilon)\langle A\rangle^{-1}.$
The term in (19) can be rewritten as
${\rm i}\langle A\rangle^{-1}(AG^{W}_{z}(\varepsilon)-G^{W}_{z}(\varepsilon)A)\langle A\rangle^{-1},$
and therefore satisfies
[…]
$\lesssim$ $1+m_{W}+\|F^{W}_{z}(\varepsilon)\|^{\frac{1}{2}}_{0}\varepsilon^{-\frac{1}{2}}.$
For the term in (19), we have
Lemma 3.3.
The commutator ${\mathcal{B}}_{W},A]$ is bounded on $\ell^{2}({\mathbb{Z}}^{d})$, with
$\left\|\left[{\mathcal{B}}_{W},A\right]\right\|_{0}\lesssim 1+M_{W}.$
Proof.
Recalling (10), we write
$\left[{\mathcal{B}}_{W},A\right]$ $=$ $[g(H_{W}),A]{\rm i}[H_{W},A]g(H_{W})+g(H_{W}){\rm i}[[H_{W},A],A]g(H_{W})$
$+\,g(H_{W}){\rm i}[H_{W},A][g(H_{W}),A].$
Arguing as in [31, Lemma 7.4], one has
$\|[g(H_{W}),A]\|_{0}\lesssim 1+m_{W}.$
Moreover, using (5) and (2), we obtain
$\|{\rm i}[H_{W},A]\|_{0}\lesssim 1+m_{W},\quad\|{\rm i}[[H_{W},A],A]\|_{0}\lesssim 1+M_{W}.$
Combining these estimates and recalling (11), we have
$\|[{\mathcal{B}}_{W},A]\|_{0}\lesssim(1+m_{W})^{2}+(1+M_{W})\lesssim 1+M_{W}.\qed$
Assume that $M_{W}>1$ and $m_{W}\lesssim 1$. The term in (19) can be estimated as
[…]
$\lesssim$ $1+M_{W}\left(\varepsilon^{\frac{1}{2}}\|F^{W}_{z}(\varepsilon)\|^{\frac{1}{2}}_{0}+\|F^{W}_{z}(\varepsilon)\|_{0}\right).$
Consequently, we obtain the differential inequality
(20) $\left\|\frac{d}{d\varepsilon}F^{W}_{z}(\varepsilon)\right\|_{0}\lesssim 1+\|F^{W}_{z}(\varepsilon)\|^{\frac{1}{2}}_{0}\left(\varepsilon^{-\frac{1}{2}}+M_{W}\varepsilon^{\frac{1}{2}}\right)+M_{W}\|F^{W}_{z}(\varepsilon)\|_{0}.$
Combining the differential inequality (20) with the a priori bound (16), one can show, following [30] or [31], that
(21) $\|F_{z}^{W}(\varepsilon)\|_{0}\lesssim M_{W}^{\frac{1}{2}}|\ln\varepsilon|^{\frac{1}{2}}\varepsilon^{\frac{1}{2}}+M_{W}^{\frac{3}{2}}|\ln\varepsilon|^{\frac{1}{2}}\varepsilon^{\frac{3}{2}}+M^{2}_{W}|\ln\varepsilon|\varepsilon.$
We begin by inserting the a priori bound (16) into (20) and integrating towards 0. This yields the improved bound
(22) $\|F^{W}_{z}(\varepsilon)\|_{0}\lesssim 1+M_{W}\varepsilon+M_{W}|\ln\varepsilon|\lesssim M_{W}|\ln\varepsilon|.$
Inserting the above bound into (20), and integrating once more, we have (21). In particular, (21) implies LAP for $H_{W}$.
In the following proposition, we show that the estimates in (14) and (15) can be improved over a prescribed range of $\varepsilon$ by means of an iterative integration argument.
Proposition 3.4.
Let $\frac{1}{2}<\sigma<1$. For $0<\varepsilon<M_{W}^{-\frac{1}{\sigma}}$, one has
(23) $\|F^{W}_{z}(\varepsilon)\|_{0}$ $\lesssim$ $\varepsilon^{2(1-\sigma)},$
(24) $\|G^{W}_{z}(\varepsilon)\langle A\rangle^{-1}\|_{0},\ \|\langle A\rangle^{-1}G^{W}_{z}(\varepsilon)\|_{0}$ $\lesssim$ $\varepsilon^{\frac{1}{2}-\sigma}.$
Proof.
For $0<\varepsilon<M_{W}^{-\frac{1}{\sigma}}$ with $\sigma<1$, we have
(25) $M_{W}\varepsilon^{\frac{1}{2}}<\varepsilon^{-\sigma+\frac{1}{2}}<\varepsilon^{-\frac{1}{2}},$
so that (20) simplifies to
(26) $\left\|\frac{d}{d\varepsilon}F^{W}_{z}(\varepsilon)\right\|_{0}\lesssim 1+\|F^{W}_{z}(\varepsilon)\|_{0}^{\frac{1}{2}}\varepsilon^{-\frac{1}{2}}+M_{W}\|F^{W}_{z}(\varepsilon)\|_{0}.$
Inserting (22) into (26), we obtain
(27) $\left\|\frac{d}{d\varepsilon}F^{W}_{z}(\varepsilon)\right\|_{0}\lesssim 1+M_{W}^{\frac{1}{2}}|\ln\varepsilon|^{\frac{1}{2}}\varepsilon^{-\frac{1}{2}}+M_{W}^{2}|\ln\varepsilon|.$
We now perform an iterative bootstrap:
[…]
We claim that for any $k\in{\mathbb{N}}^{*}$,
(28) $\left\|\frac{d}{d\varepsilon}F^{W}_{z}(\varepsilon)\right\|_{0}\lesssim_{k}1+|\ln\varepsilon|\sum_{j=0}^{k}\left(M_{W}^{j+1}\varepsilon^{j-1}\right)^{2^{-(k-j)}},$
where the implicit constant in $\lesssim_{k}$” depends on $k$. This holds for $k=1$ by (27). Assume it holds for some $k$. Integrating towards 0, we obtain
(29) $\left\|F^{W}_{z}(\varepsilon)\right\|_{0}\lesssim_{k}|\ln\varepsilon|\sum_{j=0}^{k}M_{W}^{(j+1)2^{-(k-j)}}\varepsilon^{1+(j-1)2^{-(k-j)}}.$
Inserting (29) into (26) yields
$\left\|\frac{dF^{W}_{z}}{d\varepsilon}(\varepsilon)\right\|_{0}$ $\lesssim_{k}1+|\ln\varepsilon|^{\frac{1}{2}}\sum_{j=0}^{k}\left(M_{W}^{j+1}\varepsilon^{j-1}\right)^{2^{-(k+1-j)}}$
(30) $\quad+|\ln\varepsilon|\sum_{j=0}^{k}M_{W}^{1+(j+1)2^{-(k-j)}}\varepsilon^{1+(j-1)2^{-(k-j)}}$
$\lesssim_{k}1+|\ln\varepsilon|\sum_{j=0}^{k+1}\left(M_{W}^{j+1}\varepsilon^{j-1}\right)^{2^{-(k+1-j)}},$
where the sum in (30) is estimated as
$\sum_{j=0}^{k}M_{W}^{1+(j+1)\cdot 2^{-(k-j)}}\varepsilon^{1+(j-1)\cdot 2^{-(k-j)}}$
$=$ $M_{W}^{k+2}\varepsilon^{k}+\sum_{j=0}^{k-1}\left(M_{W}^{j+2}\varepsilon^{j}\right)^{2^{-(k-j)}}\cdot(M_{W}\varepsilon)^{1-2^{-(k-j)}}$
$\leq$ $\sum_{j=1}^{k+1}\left(M_{W}^{j+1}\varepsilon^{j-1}\right)^{2^{-(k+1-j)}}.$
Thus (28) holds for all $k$.
Integrating (28) again, we recover (29). Consequently,
(31) $\|F^{W}_{z}(\varepsilon)\|^{\frac{1}{2}}_{0}\varepsilon^{-\frac{1}{2}}$ $\lesssim_{k}$ $|\ln\varepsilon|^{\frac{1}{2}}\varepsilon^{-\frac{1}{2}}\sum_{j=0}^{k}M_{W}^{(j+1)\cdot 2^{-(k+1-j)}}\varepsilon^{\frac{1}{2}+(j-1)\cdot 2^{-(k+1-j)}}$
$\leq$ $|\ln\varepsilon|^{\frac{1}{2}}\sum_{j=0}^{k}\varepsilon^{(-(j+1)\sigma+(j-1))\cdot 2^{-(k+1-j)}}.$
For $\frac{1}{2}<\sigma<1$, let us define
(32) $k=k(\sigma):=\left\lfloor\frac{1+\sigma}{1-\sigma}+\frac{1}{\ln(2)}\ln\left(\frac{1+\sigma}{2\sigma-1}\right)\right\rfloor+3,$
and for $0\leq j\leq k$,
$S_{k,j}$ $:=$ $\frac{-(j+1)\sigma+(j-1)}{2^{k+1-j}}+\left(\sigma-\frac{1}{2}\right)$
$=$ $\frac{1}{2^{k+1-j}}\left((2\sigma-1)2^{k-j}+(1-\sigma)j-(1+\sigma)\right).$
Then we have, for $\frac{1+\sigma}{1-\sigma}+1\leq j\leq k$,
$S_{k,j}>\frac{(1-\sigma)j-(1+\sigma)}{2^{k+1-j}}\geq\frac{(2\sigma-1)(1-\sigma)}{8(1+\sigma)},$
and for $0\leq j\leq\frac{1+\sigma}{1-\sigma}+1$,
$S_{k,j}>\frac{(2\sigma-1)2^{k-j}-(1+\sigma)}{2^{k+1-j}}\geq(2\sigma-1)2^{-4-\frac{1+\sigma}{1-\sigma}}.$
Both lower bounds are positive and depend only on $\sigma$. Hence,
$\|F^{W}_{z}(\varepsilon)\|^{\frac{1}{2}}_{0}\varepsilon^{-\frac{1}{2}}\cdot\varepsilon^{\sigma-\frac{1}{2}}\lesssim|\ln\varepsilon|^{\frac{1}{2}}\sum_{j=0}^{k}\varepsilon^{S_{k,j}}\lesssim 1,$
which implies (23). Using (14) and (15), we deduce (24) through
$\|G^{W}_{z}(\varepsilon)\langle A\rangle^{-1}\|_{0},\ \|\langle A\rangle^{-1}G^{W}_{z}(\varepsilon)\|_{0}\lesssim\varepsilon^{\frac{1}{2}-\sigma}.\qed$
Remark 3.5.
The iterative argument can be continued beyond $k(\sigma)$, defined in (32), yielding further improvements of the estimate for $k>k(\sigma)$. In this case, however, the implicit constants in the corresponding bounds also depend on $k$.
3.2. Proof of Theorem 3.1-(ii)
Let $V$ be a potential satisfying (2), and let $W$ be a cut-off of $V$, i.e., there exists $N\in{\mathbb{N}}$ such that
$W_{n}=\left\{\begin{array}[]{cl}V_{n},&|n|\leq N,\\ 0,&|n|>N.\end{array}\right$.
Then the quantity $m_{W}$ defined in (11) is uniformly bounded by a constant depending only on $V$. We may assume that $M_{W}\to+\infty$ as $\|V-W\|_{0}\to 0$. Indeed, if the set
$\{M_{W}:W\text{ is a cut-off of }V\}$
is bounded, then necessarily $V_{n}=O(|n|^{-2})$, and the LAP follows from Section 3.1. From now on, the implicit constants appearing in estimates denoted by $\lesssim$” and $\gtrsim$” may also depend on $V$, but are uniform with respect to all admissible cut-offs $W$.
Let $E\in[-2d,2d]\setminus{\mathcal{Z}}$. By assertion (i), the operators $H_{V}$ and $H_{W}$ have at most finitely many eigenvalues in a neighborhood of $E$ contained in $2d,2d]\setminus{\mathcal{Z}}$. In view of Lemma 2.1 and Remark 2.2, there exists a neighborhood ${\mathcal{I}}_{0}=(E-\delta_{0},E+\delta_{0})$, independent of $W$, such that
(33) $\chi_{{\mathcal{I}}_{0}}(H_{V}){\rm i}[H_{V},A]\chi_{{\mathcal{I}}_{0}}(H_{V})\geq\alpha\chi_{{\mathcal{I}}_{0}}(H_{V})+\chi_{{\mathcal{I}}_{0}}(H_{V})K_{V}\chi_{{\mathcal{I}}_{0}}(H_{V}),$
(34) $\chi_{{\mathcal{I}}_{0}}(H_{W}){\rm i}[H_{W},A]\chi_{{\mathcal{I}}_{0}}(H_{W})\geq\alpha\chi_{{\mathcal{I}}_{0}}(H_{W})+\chi_{{\mathcal{I}}_{0}}(H_{W})K_{W}\chi_{{\mathcal{I}}_{0}}(H_{W}).$
As shown in Section 3.1, for every cut-off $W$, one can obtain a strict Mourre estimate of the form (10) by suitably shrinking the interval ${\mathcal{I}}_{0}$. The following lemma shows that this procedure can be carried out uniformly with respect to $W$.
Lemma 3.6.
For $E\in[-2d,2d]\setminus{\mathcal{Z}}$, there exist
  • $0<\delta_{1}<\delta_{0}$, defining the intervals
    (35) ${\mathcal{I}}_{1}=(E-\delta_{1},E+\delta_{1}),\qquad{\mathcal{I}}_{1}^{+}=\left(E-\tfrac{\delta_{1}}{2},E+\tfrac{\delta_{1}}{2}\right),$
  • a function $g_{{\mathcal{I}}_{1}}\in C_{c}^{\infty}({\mathbb{R}},[0,1])$, supported in ${\mathcal{I}}_{1}$ and identically equal to $1$ on ${\mathcal{I}}_{1}^{+}$,
both independent of the cut-off $W$, such that
$g_{{\mathcal{I}}_{1}}(H_{V}){\rm i}[H_{V},A]g_{{\mathcal{I}}_{1}}(H_{V})$ $\geq$ $\frac{\alpha}{2}g^{2}_{{\mathcal{I}}_{1}}(H_{V}),$
$g_{{\mathcal{I}}_{1}}(H_{W}){\rm i}[H_{W},A]g_{{\mathcal{I}}_{1}}(H_{W})$ $\geq$ $\frac{\alpha}{2}g^{2}_{{\mathcal{I}}_{1}}(H_{W}).$
Proof.
We first establish a strict Mourre estimate for $H_{V}$. As in (10), the compactness of $K_{V}$ implies that there exist
  • $0<\delta_{1}^{\prime}<\frac{\delta_{0}}{2}$, defining
    ${\mathcal{I}}_{1}^{\prime}=(E-2\delta_{1}^{\prime},E+2\delta_{1}^{\prime}),\quad{\mathcal{I}}_{1}^{\prime\prime}=(E-\delta_{1}^{\prime},E+\delta_{1}^{\prime}),$
  • $g_{{\mathcal{I}}_{1}^{\prime}}\in C_{c}^{\infty}({\mathbb{R}},[0,1])$, supported in ${\mathcal{I}}_{1}^{\prime}$ and equal to $1$ on ${\mathcal{I}}_{1}^{\prime\prime}$,
such that
${\mathcal{B}}_{{\mathcal{I}}_{1}^{\prime}}^{V}:=g_{{\mathcal{I}}_{1}^{\prime}}(H_{V}){\rm i}[H_{V},A]g_{{\mathcal{I}}_{1}^{\prime}}(H_{V})\geq\frac{3\alpha}{4}g_{{\mathcal{I}}_{1}^{\prime}}^{2}(H_{V}).$
Let $W$ be a cut-off of $V$ and set X:=VWX:=V-W. Then
$\|g^{2}_{{\mathcal{I}}_{1}^{\prime}}(H_{V})-g^{2}_{{\mathcal{I}}_{1}^{\prime}}(H_{W})\|_{0}\lesssim\|X\|_{0}.$
Define
$B_{{\mathcal{I}}_{1}^{\prime}}^{X}:=g_{{\mathcal{I}}_{1}^{\prime}}(H_{V})\,{\rm i}[H_{V},A]\,g_{{\mathcal{I}}_{1}^{\prime}}(H_{V})-g_{{\mathcal{I}}_{1}^{\prime}}(H_{W})\,{\rm i}[H_{W},A]\,g_{{\mathcal{I}}_{1}^{\prime}}(H_{W}).$
A direct computation yields
(36) $\|B_{{\mathcal{I}}_{1}^{\prime}}^{X}\|_{0}\lesssim\|[X,A]\|_{0}+\|X\|_{0}.$
Under (2), $\|X\|_{0}\to 0$ implies $\|[X,A]\|_{0}\to 0$, hence there exists $\epsilon_{*}>0$ such that if $\|X\|_{0}<\epsilon_{*}$, then
$\|B_{{\mathcal{I}}_{1}^{\prime}}^{X}\|_{0}<\frac{\alpha}{8},\quad\|g^{2}_{{\mathcal{I}}_{1}^{\prime}}(H_{V})-g^{2}_{{\mathcal{I}}_{1}^{\prime}}(H_{W})\|_{0}<\frac{1}{6}.$
In this case, for ${\mathcal{B}}_{{\mathcal{I}}^{\prime}_{1}}^{W}=g_{{\mathcal{I}}^{\prime}_{1}}(H_{W}){\rm i}[H_{W},A]g_{{\mathcal{I}}^{\prime}_{1}}(H_{W})$, we have
${\mathcal{B}}_{{\mathcal{I}}^{\prime}_{1}}^{W}$ $=$ ${\mathcal{B}}_{{\mathcal{I}}^{\prime}_{1}}^{V}-B_{{\mathcal{I}}^{\prime}_{1}}^{X}$
$\geq$ $\frac{3\alpha}{4}g^{2}_{{\mathcal{I}}^{\prime}_{1}}(H_{W})+\frac{3\alpha}{4}(g^{2}_{{\mathcal{I}}^{\prime}_{1}}(H_{V})-g^{2}_{{\mathcal{I}}^{\prime}_{1}}(H_{W}))-B_{{\mathcal{I}}^{\prime}_{1}}^{X}.$
Therefore, if $0\leq\|X\|_{0}<\epsilon_{*}$, then
(37) $\chi_{{\mathcal{I}}^{\prime\prime}_{1}}(H_{W}){\rm i}[H_{W},A]\chi_{{\mathcal{I}}^{\prime\prime}_{1}}(H_{W})$ $=$ $\chi_{{\mathcal{I}}^{\prime\prime}_{1}}(H_{W}){\mathcal{B}}_{{\mathcal{I}}^{\prime}_{1}}^{W}\chi_{{\mathcal{I}}^{\prime\prime}_{1}}(H_{W})$
$\geq$ $\frac{\alpha}{2}\chi_{{\mathcal{I}}^{\prime\prime}_{1}}(H_{W}).$
On the other hand, if $\|X\|_{0}\geq\varepsilon_{*}$, then only finitely many such cut-offs $W$ arise. For each of them, the compactness of $K_{W}$ allows one to shrink the interval and obtain a strict Mourre estimate. Taking the minimum over this finite family, one can choose $0<\delta_{1}<\delta_{1}^{\prime}$ and define the intervals ${\mathcal{I}}_{1}$ and ${\mathcal{I}}_{1}^{+}$ as in (35), together with a function $g_{{\mathcal{I}}_{1}}\in C_{c}^{\infty}({\mathbb{R}},[0,1])$ supported in ${\mathcal{I}}_{1}$ and equal to $1$ on ${\mathcal{I}}_{1}^{+}$, such that
${\mathcal{B}}_{{\mathcal{I}}_{1}}^{W}=g_{{\mathcal{I}}_{1}}(H_{W}){\rm i}[H_{W},A]g_{{\mathcal{I}}_{1}}(H_{W})\geq\frac{\alpha}{2}g_{{\mathcal{I}}_{1}}^{2}(H_{W}).$
Noting that
$g_{{\mathcal{I}}_{1}}(H_{W})\chi_{{\mathcal{I}}^{\prime\prime}_{1}}(H_{W})=\chi_{{\mathcal{I}}^{\prime\prime}_{1}}(H_{W})g_{{\mathcal{I}}_{1}}(H_{W})=g_{{\mathcal{I}}_{1}}(H_{W})$
For $0\leq\|X\|_{0}<\epsilon_{*}$, the inequality (37) implies that
${\mathcal{B}}_{{\mathcal{I}}_{1}}^{W}=g_{{\mathcal{I}}_{1}}(H_{W})\chi_{{\mathcal{I}}^{\prime\prime}_{1}}(H_{W}){\rm i}[H_{W},A]\chi_{{\mathcal{I}}^{\prime\prime}_{1}}(H_{W})g_{{\mathcal{I}}_{1}}(H_{W})\\ \geq\frac{\alpha}{2}g^{2}_{{\mathcal{I}}_{1}}(H_{W}).\qed$
To prove (9) for potentials $V$ satisfying (2), we estimate the difference between the resolvents
$G^{W}_{z}(\varepsilon)=(H_{W}-z-{\rm i}\varepsilon{\mathcal{B}}_{W})^{-1},\qquad G^{V}_{z}(\varepsilon)=(H_{V}-z-{\rm i}\varepsilon{\mathcal{B}}_{V})^{-1},$
where $W$ is a suitable cut-off of $V$. Here, the positive operators
${\mathcal{B}}_{W}=g_{{\mathcal{I}}_{1}}(H_{W})\,{\rm i}[H_{W},A]\,g_{{\mathcal{I}}_{1}}(H_{W}),\quad{\mathcal{B}}_{V}=g_{{\mathcal{I}}_{1}}(H_{V})\,{\rm i}[H_{V},A]\,g_{{\mathcal{I}}_{1}}(H_{V}),$
are defined using the interval ${\mathcal{I}}_{1}$ and the function $g_{{\mathcal{I}}_{1}}\in C_{c}^{\infty}({\mathbb{R}},[0,1])$ constructed in Lemma 3.6. According to [30], for
$F^{V}_{z}(\varepsilon):=\langle A\rangle^{-1}G^{V}_{z}(\varepsilon)\langle A\rangle^{-1},$
one has
(38) $\|G^{V}_{z}(\varepsilon)\langle A\rangle^{-1}\|_{0},\ \|\langle A\rangle^{-1}G^{V}_{z}(\varepsilon)\|_{0}\lesssim 1+\varepsilon^{-\frac{1}{2}}\|F^{V}_{z}(\varepsilon)\|_{0}^{\frac{1}{2}},$
by an argument analogous to that used to prove (14) and (15). As in (36), letting X=VWX=V-W, we have
$\|B_{X}\|_{0}=\|{\mathcal{B}}_{W}-{\mathcal{B}}_{V}\|_{0}\lesssim\|[X,A]\|_{0}+\|X\|_{0}\lesssim 1.$
For $N\in{\mathbb{N}}^{*}$, define the cut-off $W_{N}$ of $V$ by
$W_{N}=(V_{n})_{|n|\leq N},\quad u_{N}:=M_{W_{N}}=\max_{|n|\leq N}\{|n|^{2}|V_{n}|\}.$
Then $u_{N}\to\infty$ as $N\to\infty$. Under the decay condition (2), one has $u_{N}=o(N)$ as $N\to\infty$. Let $N_{0}$ be the minimal index such that $u_{N_{0}}>1$. Noting that
$(u_{N_{0}},+\infty)\subset\bigcup_{N\geq N_{0}}(u_{N},10u_{N+1}),$
we deduce that, for any fixed $0<\varepsilon<u_{N_{0}}^{-\frac{1}{\sigma}}$ with $\frac{1}{2}<\sigma<1$, there exists $N\geq N_{0}$ such that
$u_{N}^{\frac{1}{\sigma}}<\varepsilon^{-1}<10^{\frac{1}{\sigma}}u_{N+1}^{\frac{1}{\sigma}}.$
We then decompose $V$ as $V=W+X$ with $W=W_{N}$.
By (23) and (24), we have
(39) $\|F_{z}^{W}(\varepsilon)\|_{0}$ $\lesssim$ $\varepsilon^{2-2\sigma},$
(40) $\|G^{W}_{z}(\varepsilon)\langle A\rangle^{-1}\|_{0},\ \|\langle A\rangle^{-1}G^{W}_{z}(\varepsilon)\|_{0}$ $\lesssim$ $\varepsilon^{\frac{1}{2}-\sigma}.$
Since $\varepsilon^{-\sigma}<10u_{N+1}$ and $u_{N+1}\lesssim N$, we have $N\gtrsim\varepsilon^{-\sigma}$. Moreover, since $V_{n}=o(|n|^{-1})$, we obtain
$\|X\|_{0}\lesssim\sup_{|n|>N}|V_{n}|\lesssim N^{-1}\lesssim\varepsilon^{\sigma}.$
Noting that
$G^{V}_{z}(\varepsilon)$ $=$ $G^{W}_{z}(\varepsilon)(H_{W}-z-{\rm i}\varepsilon{\mathcal{B}}_{W})G^{V}_{z}(\varepsilon)$
$=$ $G^{W}_{z}(\varepsilon)(1-XG^{V}_{z}(\varepsilon)+{\rm i}\varepsilon B_{X}G^{V}_{z}(\varepsilon)),$
we deduce
(41) $\|F^{V}_{z}(\varepsilon)-F^{W}_{z}(\varepsilon)\|_{0}$ $=\|\langle A\rangle^{-1}(G^{V}_{z}(\varepsilon)-G^{W}_{z}(\varepsilon))\langle A\rangle^{-1}\|_{0}$
$\leq\|\langle A\rangle^{-1}G^{W}_{z}(\varepsilon)\|_{0}\,\|X\|_{0}\,\|G^{V}_{z}(\varepsilon)\langle A\rangle^{-1}\|_{0}$
$\quad+\varepsilon\|\langle A\rangle^{-1}G^{W}_{z}(\varepsilon)\|_{0}\,\|B_{X}\|_{0}\,\|G^{V}_{z}(\varepsilon)\langle A\rangle^{-1}\|_{0}$
$\lesssim\varepsilon^{\frac{1}{2}-\sigma}(\varepsilon^{\sigma}+\varepsilon)\,\|G^{V}_{z}(\varepsilon)\langle A\rangle^{-1}\|_{0}.$
Using the bound (38), we obtain
$\|F^{V}_{z}(\varepsilon)-F^{W}_{z}(\varepsilon)\|_{0}\lesssim\|F^{V}_{z}(\varepsilon)\|_{0}^{\frac{1}{2}}.$
Combining (39) and (41), we deduce
$\|F^{V}_{z}(\varepsilon)\|_{0}\lesssim\varepsilon^{2-2\sigma}+\|F^{V}_{z}(\varepsilon)\|_{0}^{\frac{1}{2}},\qquad 0<\varepsilon<u_{N_{0}}^{-\frac{1}{\sigma}}.$
Hence $\|F_{z}^{V}(\varepsilon)\|_{0}$ is uniformly bounded, which implies (9).
4. Order1-1 Ballistic Lower Bound
The order 1-1 ballistic lower bound, i.e., the lower bound in (3) with $r=1$, is established in the following proposition. Throughout this section, the implicit constants in estimates denoted by $\lesssim$” and $\gtrsim$” are independent of both the time $t$ and the initial state $u$.
Proposition 4.1.
Given $0<\alpha<8d^{-2}$, let $I_{\alpha}\subset[-2d,2d]\setminus{\mathcal{Z}}$ be a finite union of open intervals such that there exists a compact operator $K_{V}^{\alpha}$ satisfying
(42) $\chi_{I_{\alpha}}(H_{V}){\rm i}[H_{V},A]\chi_{I_{\alpha}}(H_{V})>\alpha\chi_{I_{\alpha}}(H_{V})+\chi_{I_{\alpha}}(H_{V})K^{\alpha}_{V}\chi_{I_{\alpha}}(H_{V}).$
Let $u\in{\mathcal{H}}_{\rm ac}^{1}({\mathbb{Z}}^{d})$ with $\chi_{I_{\alpha}}(H_{V})u\neq 0$. Then
(43) $\frac{\|e^{-{\rm i}tH_{V}}u\|_{1}}{t}\gtrsim\alpha^{\frac{1}{2}}\|\chi_{I_{\alpha}}(H_{V})u\|_{0},\quad\forall\ t>0.$
Remark 4.2.
The existence of the intervals $I_{\alpha}$ such that (42) holds is guaranteed by Lemma 2.1. According to the proof of Lemma 2.1, it is enough to choose $I_{\alpha}$ so that ${\rm dist}(I_{\alpha},{\mathcal{Z}})>2^{-\frac{1}{2}}d\cdot\alpha$. Moreover, on each component of $I_{\alpha}$, the compact operator $K^{\alpha}_{V}$ can be written explicitly, and its compactness follows from the compactness of both $V$ and $[V,A]$.
Throughout the proof of Proposition 4.1, we define $B:={\rm i}[H_{V},A]$, where $A$ is the conjugate operator given by $A=-{\rm i}[H_{V},-Q^{2}]$ in the Mourre estimate. For simplicity, we omit the subscript $V$ in $H_{V}$ since the potential $V$ is fixed.
Following the arguments of [9, Proposition 1.2] and [13, Proposition 4.6], it suffices to establish a quadratic-in-time lower bound for
(44) $\|Qe^{-{\rm i}tH}u\|_{0}^{2}$ $=$ $\langle u,e^{{\rm i}tH}Q^{2}e^{-{\rm i}tH}u\rangle$
$=$ $\|Qu\|_{0}^{2}+t\langle u,Au\rangle+\int_{0}^{t}\int_{0}^{s}\left\langle e^{-{\rm i}\tau H}u,Be^{-{\rm i}\tau H}u\right\rangle\,d\tau\,ds.$
Since $u\in{\mathcal{H}}^{1}({\mathbb{Z}}^{d})$ and $A:{\mathcal{H}}^{1}({\mathbb{Z}}^{d})\to\ell^{2}({\mathbb{Z}}^{d})$ is bounded, we have
$\left|\|Qu\|_{0}^{2}+t\langle u,Au\rangle\|_{0}\right|\lesssim\|u\|_{1}^{2}+t\|u\|_{0}\|Au\|_{0}\lesssim\|u\|_{1}^{2}+t\|u\|_{0}\|u\|_{1}.$
Therefore, it suffices to prove a quadratic lower bound in $t$ for the double integral term in (44).
Let $I_{\alpha}^{c}:=[-2d,2d]\setminus I_{\alpha}$. The scalar product in the double integral in (44) can be decomposed as
$\langle e^{-{\rm i}\tau H}u,Be^{-{\rm i}\tau H}u\rangle=P_{00}(\tau)+P_{10}(\tau)+P_{01}(\tau)+P_{11}(\tau),$
with the terms defined as
$P_{00}(\tau)$ $:=$ $\left\langle e^{-{\rm i}\tau H}\chi_{I_{\alpha}}(H)u,Be^{-{\rm i}\tau H}\chi_{I_{\alpha}}(H)u\right\rangle,$
$P_{01}(\tau)$ $:=$ $\left\langle e^{-{\rm i}\tau H}\chi_{I_{\alpha}}(H)u,Be^{-{\rm i}\tau H}\chi_{I_{\alpha}^{c}}(H)u\right\rangle,$
$P_{10}(\tau)$ $:=$ $\left\langle e^{-{\rm i}\tau H}\chi_{I_{\alpha}^{c}}(H)u,Be^{-{\rm i}\tau H}\chi_{I_{\alpha}}(H)u\right\rangle,$
$P_{11}(\tau)$ $:=$ $\left\langle e^{-{\rm i}\tau H}\chi_{I_{\alpha}^{c}}(H)u,Be^{-{\rm i}\tau H}\chi_{I_{\alpha}^{c}}(H)u\right\rangle.$
The proof of Proposition 4.1 relies on the following lemmas.
Lemma 4.3.
Let $I,J\subset[-2d,2d]$ be two finite unions of intervals such that $I\cap J=\varnothing$. Then, for any $u\in{\mathcal{H}}_{\rm ac}^{0}({\mathbb{Z}}^{d})$,
(45) $\langle e^{-{\rm i}\tau H}\chi_{I}(H)u,Be^{-{\rm i}\tau H}\chi_{J}(H)u\rangle\longrightarrow 0,\qquad\tau\to\infty.$
Proof.
Let $E_{H}(\cdot)$ be the projection-valued measure of $H$ and let $\{\delta_{n}\}_{n\in{\mathbb{Z}}^{d}}$ be the standard basis of $\ell^{2}({\mathbb{Z}}^{d})$. For each $n\in{\mathbb{Z}}^{d}$, define the complex measure $\mu_{n}(\cdot):=\langle\delta_{n},E_{H}(\cdot)u\rangle$. Since $u\in{\mathcal{H}}_{\rm ac}^{0}({\mathbb{Z}}^{d})$, each $\mu_{n}$ is a.c. on $\sigma_{\rm ac}(H)$. We then have
$(e^{-{\rm i}\tau H}\chi_{I}(H)u)_{n}$ $=$ $\int_{I}e^{-{\rm i}\tau\lambda}\,d\mu_{n}(\lambda),$
$(Be^{-{\rm i}\tau H}\chi_{J}(H)u)_{n}$ $=$ $\sum_{m\in{\mathbb{Z}}^{d}}B_{nm}\int_{J}e^{-{\rm i}\tau\lambda}\,d\mu_{m}(\lambda),$
where $B_{nm}$ denote the matrix elements of $B={\rm i}[H,A]$. By (5) and (2), we have the explicit expression of $B_{nm}$.
Define the restricted measures $\mu_{n}^{I}:=\mu_{n}|_{I}$ and $\mu_{m}^{J}:=\mu_{m}|_{J}$. Then the total variation norms satisfy
$\|\mu_{n}^{I}\|_{\rm TV}=|(\chi_{I}(H)u)_{n}|,\quad\|\mu_{m}^{J}\|_{\rm TV}=|(\chi_{J}(H)u)_{m}|.$
Introduce the operator $|B|$ with matrix elements $|B|_{nm}:=|B_{nm}|$, and the $\ell^{2}$-vectors
$|\chi_{I}(H)u|:=(|(\chi_{I}(H)u)_{n}|)_{n\in{\mathbb{Z}}^{d}},\quad|\chi_{J}(H)u|:=(|(\chi_{J}(H)u)_{m}|)_{m\in{\mathbb{Z}}^{d}}.$
Then we obtain
(46) $\left|\iint_{I\times J}\sum_{n,m\in{\mathbb{Z}}^{d}}B_{nm}\,d\mu_{n}(\lambda_{1})\,d\overline{\mu_{m}}(\lambda_{2})\right|$
$\leq$ $\sum_{n,m\in{\mathbb{Z}}^{d}}|B_{nm}|\,\|\mu_{n}^{I}\|_{\rm TV}\|\mu_{m}^{J}\|_{\rm TV}$
$=$ $\sum_{n,m\in{\mathbb{Z}}^{d}}|B_{nm}||(\chi_{I}(H)u)_{n}||(\chi_{J}(H)u)_{m}|$
$=$ $\langle|\chi_{I}(H)u|,|B||\chi_{J}(H)u|\rangle$
$\leq$ $\||B|\|_{\rm op}\|\chi_{I}(H)u\|_{0}\|\chi_{J}(H)u\|_{0}.$
In particular, the above bound ensures absolute convergence. Therefore, by Fubini’s theorem,
$\left\langle e^{-{\rm i}\tau H}\chi_{I}(H)u,Be^{-{\rm i}\tau H}\chi_{J}(H)u\right\rangle$
$=$ $\sum_{n\in{\mathbb{Z}}^{d}}\sum_{m\in{\mathbb{Z}}^{d}}B_{nm}\left(\int_{I}e^{-{\rm i}\tau\lambda_{1}}\,d\mu_{n}(\lambda_{1})\right)\overline{\left(\int_{J}e^{-{\rm i}\tau\lambda_{2}}\,d\mu_{m}(\lambda_{2})\right)}$
$=$ $\iint_{I\times J}e^{-{\rm i}\tau(\lambda_{1}-\lambda_{2})}\left(\sum_{n,m\in{\mathbb{Z}}^{d}}B_{nm}\,d\mu_{n}(\lambda_{1})\,d\overline{\mu_{m}}(\lambda_{2})\right).$
Define a finite complex measure $\nu$ on $\mathbb{R}$} by
$\nu(S):=\sum_{n,m\in{\mathbb{Z}}^{d}}B_{nm}\int_{I\times J}\chi_{\{\lambda_{1}-\lambda_{2}\in S\}}(\lambda_{1},\lambda_{2})\,d\mu_{n}(\lambda_{1})\,d\overline{\mu_{m}}(\lambda_{2}),\quad S\subset{\mathbb{R}}.$
Then, by (46), $|\nu|({\mathbb{R}})<\infty$, and we obtain
$\langle e^{-{\rm i}\tau H}\chi_{I}(H)u,\,Be^{-{\rm i}\tau H}\chi_{J}(H)u\rangle=\int_{{\mathbb{R}}}e^{-{\rm i}\tau E}\,d\nu(E).$
Since $I\cap J=\varnothing$, the support of $\nu$ does not contain 0. By the Riemann–Lebesgue lemma, it follows that
$\int_{{\mathbb{R}}}e^{-{\rm i}\tau E}\,d\nu(E)\longrightarrow 0,\qquad\tau\to\infty,$
which proves (45).∎
Lemma 4.4.
Given $I\subset[-2d,2d]$ a finite union of intervals with
(47) $\chi_{I}(H)B\chi_{I}(H)\geq a\chi_{I}(H)+K,$
for some $a>0$ and a compact operator $K$, we have, for $u\in{\mathcal{H}}_{\rm ac}^{0}({\mathbb{Z}}^{d})$, for $\tau>0$ sufficiently large,
(48) $\left\langle e^{-{\rm i}\tau H}\chi_{I}(H)u,Be^{-{\rm i}\tau H}\chi_{I}(H)u\right\rangle\geq\frac{a}{4}\|\chi_{I}(H)u\|_{0}^{2}.$
Proof.
Similar to the proof of (10), we can use the compactness of $C$ to obtain a strict Mourre estimate based on (47) by shrinking $I$. For $E\in I\setminus\sigma_{\rm pp}\left(H\right)$, there exists
  • $\delta>0$ sufficiently small, through which we define
    ${\mathcal{I}}^{\prime}=(E-2\delta,E+2\delta)\subset I\setminus\sigma_{\rm pp}(H),\quad{\mathcal{I}}=(E-\delta,E+\delta),$
  • $g\in C_{c}^{\infty}({\mathbb{R}},[0,1])$, supported on $\mathcal{I}$$^{\prime}$ and identically $1$ on $\mathcal{I}$},
such that
$\pm g(H)Kg(H)\leq\frac{\alpha}{2}g^{2}(H),\quad g(H){\rm i}[H,A]g(H)>\frac{\alpha}{2}g^{2}(H),$
which implies
(49) $\chi_{{\mathcal{I}}}(H){\rm i}[H,A]\chi_{{\mathcal{I}}}(H)>\frac{\alpha}{2}\chi_{{\mathcal{I}}}(H).$
Let us choose finitely many pairwise disjoint intervals
${\mathcal{I}}_{j}\subset I\setminus\sigma_{\rm pp}(H),\quad 1\leq j\leq N,$
such that the estimate (49) holds for each ${\mathcal{I}}_{j}=(E_{j}-\delta,E_{j}+\delta)$, and
(50) $\left\|\chi_{I}(H)u-\sum_{1\leq j\leq N}\chi_{{\mathcal{I}}_{j}}(H)u\right\|_{0}\leq\epsilon,$
for some arbitrarily small $\epsilon>0$. Then we can write
(52) $\left\langle e^{-{\rm i}\tau H}\chi_{I}(H)u,Be^{-{\rm i}\tau H}\chi_{I}(H)u\right\rangle$
$=$ $\sum_{1\leq j\leq N}\left\langle e^{-{\rm i}\tau H}\chi_{{\mathcal{I}}_{j}}\left(H\right)u,Be^{-{\rm i}\tau H}\chi_{{\mathcal{I}}_{j}}\left(H\right)u\right\rangle$
$+\,\sum_{1\leq j,l\leq N\atop{j\neq l}}\left\langle e^{-{\rm i}\tau H}\chi_{{\mathcal{I}}_{j}}\left(H\right)u,Be^{-{\rm i}\tau H}\chi_{{\mathcal{I}}_{l}}\left(H\right)u\right\rangle+{\mathcal{O}}(\epsilon),$
where the remainder term ${\mathcal{O}}(\epsilon)$ is uniform in $\tau$. By Lemma 4.3, the sum in (52) converges to 0 as $\tau\to\infty$. On the other hand, applying (49) with ${\mathcal{I}}={\mathcal{I}}_{j}$, we have
$\sum_{1\leq j\leq N}\left\langle e^{-{\rm i}\tau H}\chi_{{\mathcal{I}}_{j}}\left(H\right)u,Be^{-{\rm i}\tau H}\chi_{{\mathcal{I}}_{j}}\left(H\right)u\right\rangle$ $\geq$ $\frac{a}{2}\sum_{1\leq j\leq N}\|\chi_{{\mathcal{I}}_{j}}\left(H\right)u\|^{2}_{0}$
$\geq$ $\frac{a}{2}\|\chi_{I}\left(H\right)u\|^{2}_{0}-{\mathcal{O}}(\epsilon).$
Since $\epsilon>0$ is arbitrarily small, the lower bound (48) follows.∎
Proof of Proposition 4.1. By Lemma 4.3, we have
(53) $P_{01}(\tau),\ P_{10}(\tau)\to 0,\quad\tau\to\infty.$
Moreover, applying Lemma 4.4 together with the Mourre estimate (42), we obtain that, for $\tau>0$ sufficiently large,
(54) $P_{00}(\tau)\geq\frac{\alpha}{4}\|\chi_{I_{\alpha}}(H)u\|_{0}^{2}.$
For $0<\epsilon<\alpha$ sufficiently small, let $I_{\epsilon}$ be the union of intervals such that, for a compact operator $K^{\epsilon}_{V}$,
(55) $\chi_{I_{\epsilon}}(H){\rm i}[H,A]\chi_{I_{\epsilon}}(H)>\epsilon\chi_{I_{\epsilon}}(H)+\chi_{I_{\epsilon}}(H)K^{\epsilon}_{V}\chi_{I_{\epsilon}}(H).$
Since $I_{\alpha}^{c}=(I_{\epsilon}\setminus I_{\alpha})\cup I_{\epsilon}^{c}$, we decompose $P_{11}(\tau)$ as
(59) $P_{11}(\tau)$ $=$ $\left\langle e^{-{\rm i}\tau H}\chi_{I_{\epsilon}\setminus I_{\alpha}}(H)u,Be^{-{\rm i}\tau H}\chi_{I_{\epsilon}\setminus I_{\alpha}}(H)u\right\rangle$
$+\,\left\langle e^{-{\rm i}\tau H}\chi_{I_{\epsilon}\setminus I_{\alpha}}(H)u,Be^{-{\rm i}\tau H}\chi_{I_{\epsilon}^{c}}(H)u\right\rangle$
$+\,\left\langle e^{-{\rm i}\tau H}\chi_{I_{\epsilon}^{c}}(H)u,Be^{-{\rm i}\tau H}\chi_{I_{\epsilon}\setminus I_{\alpha}}(H)u\right\rangle$
$+\,\left\langle e^{-{\rm i}\tau H}\chi_{I_{\epsilon}^{c}}(H)u,Be^{-{\rm i}\tau H}\chi_{I_{\epsilon}^{c}}(H)u\right\rangle.$
By Lemma 4.4, the Mourre estimate on $I_{\epsilon}\setminus I_{\alpha}\subset I_{\epsilon}$ implies that the term (59) is non-negative. By Lemma 4.3, the cross terms (59) and (59) tend to 0 as $\tau\to\infty$. In view of Remark 4.2, we have
$I_{\epsilon}^{c}=[-2d,2d]\setminus I_{\epsilon}\subset\{E\in[-2d,2d]:{\rm dist}(E,{\mathcal{Z}})\leq\epsilon\}.$
Then $\chi_{I_{\epsilon}^{c}}(H)$ converges to 0 strongly as $\epsilon\to 0$. In particular, for $\epsilon$ sufficietly small,
$\|\chi_{I_{\epsilon}^{c}}(H)u\|_{0}\leq\frac{\alpha^{\frac{1}{2}}}{5(1+\|B\|)^{\frac{1}{2}}}\|\chi_{I_{\alpha}}(H)u\|_{0},$
and consequently, the term (59) is bounded by $\frac{\alpha}{20}\|\chi_{I_{\alpha}}(H)u\|^{2}_{0}$. Therefore, we conclude that for $\tau$ large enough,
(60) $P_{11}(\tau)\geq-\frac{\alpha}{10}\|\chi_{I_{\alpha}}(H)u\|_{0}^{2}.$
Combining (53), (54) and (60), we obtain, for $\tau$ sufficiently large,
$\langle e^{-{\rm i}\tau H}u,Be^{-{\rm i}\tau H}u\rangle\geq\frac{\alpha}{8}\|\chi_{I_{\alpha}}(H)u\|_{0}^{2}.$
Consequently, the double integral in (44) exhibits a $t^{2}$-growth, which yields the ballistic lower bound (43).∎
5. Proof of Theorem 1.2
In this section, we comlete the proof of Theorem 1.2. We will split it into two parts, presented in separate subsections, one addressing the upper bound and one addressing the lower bound in (3).
5.1. Ballistic Upper Bound
By general principles, the upper bound in (3) holds for every $r>0$ and for any self-adjoint discrete Schrödinger operator. Throughout this subsection, we only assume that $V$ is real-valued, that is, $V:{\mathbb{Z}}^{d}\to{\mathbb{R}}$, and we do not require that $V$ satisfy the decay condition (2).
Such a result is well known; compare, for example, the discussions in [1, Appendix B], [6, Lemma B.2], [7, Theorem A.1], [10, Theorem 2.6.2], [14, Theorem 2.22], and [37, Theorem 3]. [^3] There is also a ballistic upper bound result in the continuum case due to Radin and Simon [32]. For the reader’s convenience, we include a proof below. In fact, in doing so we establish a particular instance of the upper bound, (62), that is crucial in our discussion of the orderr-r ballistic lower bound for $0<r<1$ below; compare Lemma 5.4.
Proposition 5.1.
For any self-adjoint discrete Schrödinger operator $H_{V}=-\Delta+V$ on $\ell^{2}({\mathbb{Z}}^{d})$, we have the ballistic upper bound: there exists a constant $c_{r}>0$, independent of $u$, such that
(61) $\|e^{-{\rm i}tH_{V}}u\|_{r}\leq c_{r}\|u\|_{r}t^{r},\quad\forall\ u\in{\mathcal{H}}^{r}({\mathbb{Z}}^{d}),\quad t\geq 1.$
Moreover, there exists $\tilde{c}_{2}>0$ such that
(62) $\|e^{-{\rm i}tH_{V}}u\|_{2}\leq\|u\|_{2}+\|u\|_{1}t+\tilde{c}_{2}\|u\|_{0}t^{2},\quad\forall\ u\in{\mathcal{H}}^{2}({\mathbb{Z}}^{d}),\;t\geq 1.$
Proof.
Given $m\in{\mathbb{N}}^{*}$, for $u\in{\mathcal{H}}^{m}({\mathbb{Z}}^{d})$, we have
(63) $e^{{\rm i}tH_{V}}Q^{m}e^{-{\rm i}tH_{V}}u=Q^{m}u+{\rm i}\int_{0}^{t}e^{{\rm i}sH_{V}}[H_{V},Q^{m}]e^{-{\rm i}sH_{V}}u\,ds,$
where $H_{V},Q^{m}]$ is well-defined on ${\mathcal{H}}^{m-1}({\mathbb{Z}}^{d})$. Indeed, noting that $H_{V},Q^{m}]=[Q^{m},\Delta]$ for $H_{V}$ with any potential $V$, we have, for $\psi\in{\mathcal{H}}^{m-1}({\mathbb{Z}}^{d})$ and $n\in{\mathbb{Z}}^{d}$,
$([H_{V},Q^{m}]\psi)_{n}$ $=$ $([Q^{m},\Delta]\psi)_{n}$
$=$ $\sum_{j=1}^{d}\left((1+|n|^{2})^{\frac{m}{2}}-(1+|n+e_{j}|^{2})^{\frac{m}{2}}\right)\psi_{n+e_{j}}$
$+\,\sum_{j=1}^{d}\left((1+|n|^{2})^{\frac{m}{2}}-(1+|n-e_{j}|^{2})^{\frac{m}{2}}\right)\psi_{n-e_{j}},$
which implies that, there is some constant $\tilde{c}_{m}>0$, such that
$\|[H_{V},Q^{m}]\psi\|_{0}\leq 2\tilde{c}_{m}\|Q^{m-1}\psi\|_{0}=2\tilde{c}_{m}\|\psi\|_{m-1},$
since, for $j=1,\cdots,d$, for $n\in{\mathbb{Z}}^{d}$,
$\left|(1+|n|^{2})^{\frac{m}{2}}-(1+|n\pm e_{j}|^{2})^{\frac{m}{2}}\right|$ $=$ $(1+|n|^{2})^{\frac{m}{2}}\left|\left(1+\frac{2n_{j}\pm 1}{1+|n|^{2}}\right)^{\frac{m}{2}}-1\right|$
$\leq$ $\frac{\tilde{c}_{m}}{d}(1+|n|^{2})^{\frac{m-1}{2}}.$
Hence, when $V$ is a real potential, we have, through (63),
$\|e^{-{\rm i}tH_{V}}u\|_{m}$ $=$ $\|e^{{\rm i}tH_{V}}Q^{m}e^{-{\rm i}tH_{V}}u\|_{0}$
$\leq$ $\|Q^{m}u\|_{0}+\int_{0}^{t}\|[H_{V},Q^{m}]e^{-{\rm i}sH_{V}}u\|_{0}\,ds$
$\leq$ $\|u\|_{m}+2\tilde{c}_{m}\int_{0}^{t}\|e^{-{\rm i}sH_{V}}u\|_{m-1}\,ds,\quad\forall\ t\geq 0.$
Since for $m=1$, $\|e^{-{\rm i}tH_{V}}u\|_{m-1}=\|u\|_{0}$, we obtain the linear upper bound for $\|e^{-{\rm i}tH_{V}}u\|_{1}$ for $u\in{\mathcal{H}}^{1}({\mathbb{Z}}^{d})$:
$\|e^{-{\rm i}tH_{V}}u\|_{1}\leq\|u\|_{1}+2\tilde{c}_{2}\|u\|_{0}t,\quad\forall\ t\geq 0.$
Then, via induction, for $u\in{\mathcal{H}}^{2}({\mathbb{Z}}^{d})$,
$\|e^{-{\rm i}tH_{V}}u\|_{2}$ $\leq$ $\|u\|_{2}+\int_{0}^{t}\|e^{-{\rm i}sH_{V}}u\|_{1}\,ds$
$\leq$ $\|u\|_{2}+\|u\|_{1}t+\tilde{c}_{2}\|u\|_{0}t^{2},\quad\forall\ t\geq 0.$
which shows (62). The above inequalities also imply that, for $m=1,2$, there exists $c^{\prime}_{m}>0$ such that
$\|e^{-{\rm i}tH_{V}}u\|_{m}\leq c^{\prime}_{m}\|u\|_{m}(1+t)^{m},\quad\forall\ t\geq 0,$
through which we obtain (61) for $r=1,2$.
Assume that, for $m\in{\mathbb{N}}^{*}$, there exists a constant $c^{\prime}_{m}>0$ such that, for any $u\in{\mathcal{H}}^{m}({\mathbb{Z}}^{d})$,
$\|e^{-{\rm i}tH_{V}}u\|_{m}\leq c^{\prime}_{m}\|u\|_{m}(1+t)^{m},\quad\forall\ t\geq 0.$
By induction, we have
$\|e^{-{\rm i}tH_{V}}u\|_{m+1}$ $\leq$ $\|u\|_{m+1}+2\tilde{c}_{m+1}\int_{0}^{t}\|e^{-{\rm i}sH_{V}}u\|_{m}\,ds$
$\leq$ $\|u\|_{m+1}+\frac{2\tilde{c}_{m+1}c^{\prime}_{m}}{m+1}\|u\|_{m}(1+t)^{m+1},\quad\forall\ t\geq 0.$
Therefore, the upper bound (61) is shown when $r>0$ is an integer.
For $r\in(m,m+1)$ with some $m\in{\mathbb{N}}$ ($\{0,1,2,\ldots\}$) we have
$r=\left((m+1)-r\right)m+\left(r-m\right)(m+1).$
Applying Hölder’s inequality, we obtain, for $\psi\in{\mathcal{H}}^{m+1}({\mathbb{Z}}^{d})$,
(64) $\|\psi\|^{2}_{r}$ $=$ $\sum_{n\in{\mathbb{Z}}^{d}}(1+|n|^{2})^{\left((m+1)-r\right)m+\left(r-m\right)(m+1)}\left|\psi_{n}\right|^{2}$
$\leq$ $\left(\sum_{n\in{\mathbb{Z}}^{d}}(1+|n|^{2})^{m}\left|\psi_{n}\right|^{2}\right)^{(m+1)-r}\left(\sum_{n\in{\mathbb{Z}}^{d}}(1+|n|^{2})^{m+1}\left|\psi_{n}\right|^{2}\right)^{r-m}$
$=$ $\|\psi\|_{m}^{2(m+1-r)}\|\psi\|_{m+1}^{2(r-m)}.$
In view of the integer order ballistic upper bound, for $u\in{\mathcal{H}}^{m+1}({\mathbb{Z}}^{d})$,
(65) $\|e^{-{\rm i}tH_{V}}u\|_{r}$ $\leq$ $\|e^{-{\rm i}tH_{V}}u\|_{m}^{m+1-r}\|e^{-{\rm i}tH_{V}}u\|_{m+1}^{r-m}$
$\leq$ $c^{m+1-r}_{m}c^{r-m}_{m+1}\|u\|_{m}^{m+1-r}\|u\|_{m+1}^{r-m}\cdot t^{\left((m+1)-r\right)m+\left(r-m\right)(m+1)}$
$\leq$ $c_{r}\|u\|_{r}t^{r},\quad{\rm with}\ c_{r}:=c^{m+1-r}_{m}c^{r-m}_{m+1}.$
Note that ${\mathcal{H}}^{m+1}({\mathbb{Z}}^{d})$ is dense in ${\mathcal{H}}^{r}({\mathbb{Z}}^{d})$ in the sense that, if $u\in{\mathcal{H}}^{r}({\mathbb{Z}}^{d})$, then for $\varepsilon>0$, there exists $u_{\varepsilon}\in{\mathcal{H}}^{m+1}({\mathbb{Z}}^{d})$ such that $\|u_{\varepsilon}-u\|_{r}<\varepsilon$. Since (65) implies that, for fixed $t\geq 0$, the time-evolution $e^{-{\rm i}tH_{V}}:{\mathcal{H}}^{m+1}({\mathbb{Z}}^{d})\to{\mathcal{H}}^{r}({\mathbb{Z}}^{d})$ is bounded with operator norm controlled by $c_{r}t^{r}$. By density and the continuous extension theorem for bounded linear operators defined on a dense subspace, $e^{-{\rm i}tH_{V}}$ extends uniquely to a bounded operator on ${\mathcal{H}}^{r}({\mathbb{Z}}^{d})$ and (61) is satisfied.∎
5.2. Ballistic Lower Bound
Let us next focus on establishing the lower bound in (3) for $H_{V}$ with the potential $V$ satisfying (2). For $r=1$, it is shown in Proposition 4.1. The orderr-r ballistic lower bound for $r\geq 1$ is deduced from the following lemma.
Lemma 5.2.
For any self-adjoint operator $H$ on $\ell^{2}({\mathbb{Z}}^{d})$, if there exists $r>0$ such that for any non-vanishing $u\in{\mathcal{H}}_{\rm ac}^{r}({\mathbb{Z}}^{d})$, $\|e^{-{\rm i}tH}u\|_{r}\geq b_{r,u}t^{r}$ for some constant $b_{r,u}>0$, then for any $r^{\prime}>r$ and $u\in{\mathcal{H}}_{\rm ac}^{r^{\prime}}({\mathbb{Z}}^{d})\setminus\{0\}$, there exists a constant $b_{r^{\prime},u}>0$ such that $\|e^{-{\rm i}tH}u\|_{r^{\prime}}\geq b_{r^{\prime},u}t^{r^{\prime}}$.
Remark 5.3.
This argument is a variation of the argument given in the proof of [14, Lemma 2.7].
Proof.
Suppose that $u\in{\mathcal{H}}_{\rm ac}^{r}({\mathbb{Z}}^{d})$ is $\ell^{2}$-normalized. By assumption, there exists $b_{r,u}>0$ such that $\|e^{-{\rm i}tH}u\|_{r}\geq b_{r,u}t^{r}$, which means that
$\sum_{n\in{\mathbb{Z}}^{d}}(1+|n|^{2})^{r}|\langle\delta_{n},e^{-{\rm i}tH}u\rangle|^{2}\geq b^{2}_{r,u}t^{2r}.$
For $r^{\prime}>r$, note that $x\mapsto|x|^{\frac{r^{\prime}}{r}}$ is convex on $\mathbb{R}$}. Using Jensen’s inequality (and emphasizing that $u$ is $\ell^{2}$-normalized so that $n\mapsto|\langle\delta_{n},e^{-{\rm i}tH}u\rangle|^{2}$ is a probability distribution on ${\mathbb{Z}}^{d}$), we observe that, for any $\ell^{2}$-normalized $u\in{\mathcal{H}}_{\rm ac}^{r^{\prime}}({\mathbb{Z}}^{d})\subset{\mathcal{H}}_{\rm ac}^{r}({\mathbb{Z}}^{d})$,
$\sum_{n\in{\mathbb{Z}}^{d}}(1+|n|^{2})^{r^{\prime}}|\langle\delta_{n},e^{-{\rm i}tH}u\rangle|^{2}$ $=\sum_{n\in{\mathbb{Z}}^{d}}\left((1+|n|^{2})^{r}\right)^{\frac{r^{\prime}}{r}}|\langle\delta_{n},e^{-{\rm i}tH}u\rangle|^{2}$
$\geq\left(\sum_{n\in{\mathbb{Z}}^{d}}(1+|n|^{2})^{r}|\langle\delta_{n},e^{-{\rm i}tH}u\rangle|^{2}\right)^{\frac{r^{\prime}}{r}}$
$\geq b^{\frac{2r^{\prime}}{r}}_{r,u}t^{2r^{\prime}},$
which implies that $\|e^{-{\rm i}tH}u\|_{r}\geq b_{r^{\prime},u}t^{r^{\prime}}$ with $b_{r^{\prime},u}=b_{r,u}^{\frac{r^{\prime}}{r}}$. By scaling, the bound extends to non-normalized $u\in{\mathcal{H}}_{\rm ac}^{r^{\prime}}({\mathbb{Z}}^{d})\setminus\{0\}$.∎
As for the case $0<r<1$, we have the following lemma, which completes the proof of Theorem 1.2.
Lemma 5.4.
For the Schrödinger operator $H_{V}$ with $V$ satisfying (2), for any $0<r<1$ and any non-vanishing $u\in{\mathcal{H}}_{\rm ac}^{r}({\mathbb{Z}}^{d})$, there exists a constant $b_{r,u}>0$ such that for $t\geq 1$,
$\|e^{-{\rm i}tH_{V}}u\|_{r}\geq b_{r,u}t^{r}.$
Proof.
Applying Hölder’s inequality as (64), we obtain, for $0<r<1$,
$\|e^{-{\rm i}tH_{V}}u\|_{1}\leq\|e^{-{\rm i}tH_{V}}u\|_{r}^{\frac{1}{2-r}}\|e^{-{\rm i}tH_{V}}u\|^{\frac{1-r}{2-r}}_{2}.$
According to (62) in Proposition 5.1 and Proposition 4.1, for any $u\in{\mathcal{H}}^{2}_{\rm ac}({\mathbb{Z}}^{d})\setminus\{0\}$, one can find $0<\alpha<2$ and $I_{\alpha}\subset[-2d,2d]\setminus{\mathcal{Z}}$ such that $\chi_{I_{\alpha}}(H_{V})u$ is non-vanishing and there exists $a_{1}>0$ such that
$\|e^{-{\rm i}tH_{V}}u\|_{1}$ $\geq$ $a_{1}\alpha^{\frac{1}{2}}\|\chi_{I_{\alpha}}(H_{V})u\|_{0}\,t,$
$\|e^{-{\rm i}tH_{V}}u\|_{2}$ $\leq$ $\|u\|_{2}+\|u\|_{1}\,t+\tilde{c}_{2}\|u\|_{0}\,t^{2},$
Hence, as $t\to\infty$,
$\|e^{-{\rm i}tH_{V}}u\|_{r}\geq\frac{(a_{1}\alpha^{\frac{1}{2}}\|\chi_{I_{\alpha}}(H_{V})u\|_{0}t)^{2-r}}{(\|u\|_{2}+\|u\|_{1}t+\tilde{c}_{2}\|u\|_{0}t^{2})^{1-r}}\geq\frac{(a_{1}\alpha^{\frac{1}{2}}\|\chi_{I_{\alpha}}(H_{V})u\|_{0})^{2-r}}{2^{1-r}(\tilde{c}_{2}\|u\|_{0})^{1-r}}t^{r}.$
Note that ${\mathcal{H}}^{2}_{\rm ac}({\mathbb{Z}}^{d})={\mathcal{H}}^{2}({\mathbb{Z}}^{d})\cap{\mathcal{H}}_{\rm ac}(H_{V})$ is dense in ${\mathcal{H}}^{r}_{\rm ac}({\mathbb{Z}}^{d})={\mathcal{H}}^{r}({\mathbb{Z}}^{d})\cap{\mathcal{H}}_{\rm ac}(H_{V})$ in the sense that, given $u\in{\mathcal{H}}_{\rm ac}^{r}({\mathbb{Z}}^{d})$, for any $\varepsilon>0$, there exists $u_{\varepsilon}\in{\mathcal{H}}_{\rm ac}^{2}({\mathbb{Z}}^{d})$ such that $\|u_{\varepsilon}-u\|_{r}<\varepsilon$, which implies that, for $\varepsilon$ sufficiently small,
$\|\chi_{I_{\alpha}}(H_{V})u_{\varepsilon}\|_{0}$ $\geq$ $\|\chi_{I_{\alpha}}(H_{V})u\|_{0}-\|u_{\varepsilon}-u\|_{0}$
$\geq$ $\|\chi_{I_{\alpha}}(H_{V})u\|_{0}-\|u_{\varepsilon}-u\|_{r}\ \geq\ \frac{1}{2}\|\chi_{I_{\alpha}}(H_{V})u\|_{0},$
$\|u_{\varepsilon}\|_{0}$ $\leq$ $\|u\|_{0}+\|u_{\varepsilon}-u\|_{0}\ \leq\ 2\|u\|_{0}.$
According to (61) in Proposition 5.1, for $t\geq 0$,
$\|e^{-{\rm i}tH_{V}}(u_{\varepsilon}-u)\|_{r}\leq c_{r}\|u_{\varepsilon}-u\|_{r}t^{r}.$
Then, for $\varepsilon$ sufficiently small, as $t\to\infty$, we have
$\|e^{-{\rm i}tH_{V}}u\|_{r}$ $\geq\|e^{-{\rm i}tH_{V}}u_{\varepsilon}\|_{r}-\|e^{-{\rm i}tH_{V}}(u_{\varepsilon}-u)\|_{r}$
$\geq\frac{(a_{1}\alpha^{\frac{1}{2}}\|\chi_{I_{\alpha}}(H_{V})u_{\varepsilon}\|_{0})^{2-r}}{2^{1-r}(\tilde{c}_{2}\|u_{\varepsilon}\|_{0})^{1-r}}\cdot t^{r}-c_{r}\|u_{\varepsilon}-u\|_{r}t^{r}$
$\geq\left(\frac{(a_{1}\alpha^{\frac{1}{2}}\|\chi_{I_{\alpha}}(H_{V})u\|_{0})^{2-r}}{2^{4-3r}(\tilde{c}_{2}\|u\|_{0})^{1-r}}-c_{r}\varepsilon\right)t^{r}$
$\geq\frac{(a_{1}\alpha^{\frac{1}{2}}\|\chi_{I_{\alpha}}(H_{V})u\|_{0})^{2-r}}{16(\tilde{c}_{2}\|u\|_{0})^{1-r}}t^{r}.\qed$
Acknowledgment
We are grateful to Sergey Denisov, Jake Fillman, and Tal Malinovitch for helpful information about the literature and to Marc-Adrien Mandich for encouraging us to clarify in detail the absence of singular continuous spectrum assuming only (2).